### v1

# VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization

Abstract
Unprecedented breakthroughs in Large Language Models (LLMs) has amplified its penetration into application of automated visualization code generation. Few-shot prompting and query expansion techniques have notably enhanced data visualization performance, however, still fail to overcome ambiguity and complexity of natural language queries - imposing an inherent burden for manual human intervention. To mitigate such limitations, we propose a holistic framework VisPath : A Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation, which systematically enhances code quality through structured reasoning and refinement. VisPath is a multi-stage framework, specially designed to handle underspecified queries. To generate a robust final visualization code, it first utilizes initial query to generate diverse reformulated queries via Chain-of-Thought (CoT) prompting, each representing a distinct reasoning path. Refined queries are used to produce candidate visualization scripts, consequently executed to generate multiple images. Comprehensively assessing correctness and quality of outputs, VisPath generates feedback for each image, which are then fed to aggregation module to generate optimal result. Extensive experiments on benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath significantly outperforms state-of-the-art (SOTA) methods, increased up to average 17%, offering a more reliable solution for AI-driven visualization code generation.
VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
Wonduk Seo1thanks: Equal contribution.   Seungyong Lee2*   Daye Kang1,3   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author. 1Enhans.ai   2Unaffiliated   3KAIST   4Tsinghua University 1{seowonduk,seunghyun,daye}@enhans.ai 2leesy7197@khu.ac.kr   4yzh23@tsinghua.edu.cn
1 Introduction
Figure 1: Overview of different approaches for visualization code generation. Comparing two baseline methods, namely Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024), with our proposed VisPath framework.
Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets Vondrick et al. (2013); Demiralp et al. (2017); Unwin (2020); Li et al. (2024a). Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js Barrett et al. (2005); Bisong and Bisong (2019); Zhu (2013). This approach demands programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al. (2018); Sharif et al. (2024). As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible Wang et al. (2015); Dibia and Demiralp (2019); Qian et al. (2021).
In response to this challenge, Large Language Models (LLMs) have emerged as a promising solution for simplifying visualization creation Wang et al. (2023a); Han et al. (2023); Xie et al. (2024). By translating natural language instructions into executable code, LLM-based systems eliminate the need for extensive programming knowledge, allowing users to generate visualizations more intuitively Xiao et al. (2023); Ge et al. (2023); Zhang et al. (2024b). Recent visualization methods such as ChartGPT Tian et al. (2024) and NL4DV Sah et al. (2024) demonstrate the potential of LLMs to provide interactive, conversational interfaces for visualization. These systems enable users to create complex charts with minimal effort, bridging the gap between technical expertise and effective data exploration Dibia (2023); Kim et al. (2024).
More recently, LLM-based visualization frameworks such as Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al. (2024) have been introduced to improve automated visualization code generation. Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation. However, these methods face several limitations: ① both generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; ② both rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries. ③ both approaches encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback. Without a mechanism to retrieve outputs that reflect diverse possibilities, they struggle to capture intricate details, ultimately limiting the precision and adaptability of the generated visualizations.
To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a transformative approach that redefines how visualization code is generated. Traditional single-pass methods may seem efficient, but they often fall short of delivering the depth and precision users truly need. They generate code quickly, but struggle to capture the intricate details that make a visualization not just functional but meaningful. VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
Rather than simply translating user input into code, VisPath ensures that every critical aspect—both explicitly stated and implicitly necessary—is carefully considered, creating a visualization that is not only correct but insightful. At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting. These multiple candidates are evaluated using a Vision-Language Model (VLM) to assess accuracy, clarity, and alignment with the intended message. The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact. By shifting from a single-pass, one-size-fits-all approach to a dynamic, multi-layered process, VisPath sets a new standard for visualization generation—one that is not just about creating code, but about ensuring data is represented in its most clear, meaningful, and compelling form.
Extensive experiments on benchmark datasets, including MatPlotBench Yang et al. (2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate VisPath’s superiority over the state-of-the-art (SOTA) visualization generation methods. By systematically generating and evaluating multiple reasoning paths and leveraging iterative feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents. Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
2 Related Work
Numerous methods have been applied for Text-to-Visualization (Text2Vis) generation, which has significantly evolved over the years, adapting to new paradigms in data visualization and natural language processing Dibia and Demiralp (2019); Wu et al. (2022); Chen et al. (2022a, b); Rashid et al. (2022); Zhang et al. (2024a). Early approaches such as Voyager Wongsuphasawat et al. (2015) and Eviza Setlur et al. (2016) largely relied on rule-based systems, which mapped textual commands to predefined chart templates or specifications through handcrafted heuristics de Araújo Lima and Diniz Junqueira Barbosa (2020). While these methods demonstrated the feasibility of automatically converting text into visualizations Moritz et al. (2018); Cui et al. (2019), they often required extensive domain knowledge and struggled with more nuanced or ambiguous user requirements Li et al. (2021); Wang et al. (2023b). Inspired by developments in deep learning, researchers began to incorporate neural networks to handle free-form natural language and broaden the range of supported visualization types Liu et al. (2021); Luo et al. (2021).
Figure 2: Overview of the proposed VisPath framework for creating robust visualization code generation. The framework consists of combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
Building on these machine learning strategies, numerous studies have utilized Large Language Models (LLMs) to further enhance system flexibility. Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al. (2024b) utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting. More recent approaches, such as MatPlotAgent Yang et al. (2024) and PlotGen Goswami et al. (2025), extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations. The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks. By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments. This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification—ironically undermining the intended efficiency of automation. To address these limitations, we introduce VisPath, a novel framework that integrates multi-path reasoning with feedback from Vision-Language Models (VLMs) to enhance visualization code generation. Compared to conventional methods constrained by a single reasoning path, VisPath dynamically explores multiple solution pathways, improving interpretability, minimizing manual intervention, and adapting more effectively to diverse visualization requirements.
3 Methodology
We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback. VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a Vision-Language Model (VLM) evaluates and refines the outputs to ensure generation robust visualization code. An overview of this process is shown in Figure 2.
3.1 Multi-Path Generation
One of the biggest pitfalls in visualization code generation is rigid interpretation, a single query can have multiple valid visual representations depending on its dataset structure. VisPath mitigates this by generating multiple extended queries within a single interaction. Given a user query $Q$ and a corresponding dataset description $D$, a Multi-Path Agent is employed to expand the query into $K$ distinct reasoning pathways:
$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}=f_{\_}{\text{mpa}}(Q,D),$ (1)
where $f_{\_}{\text{mpa}}$ denotes the function of the Multi-Path Agent implemented via an LLM. The dataset description $D$ plays a crucial core in shaping these reasoning paths by providing contextual information about variable types, inherent relationships, and the suitability of different chart types for a more grounded interpretation of the query. Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request. This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent—even when queries are ambiguous or underspecified.
3.2 Code Generation from Reasoning Paths
Once diverse reasoning paths are established, the next stage involves translating each path into executable Python scripts. For every reasoning path $R_{\_}i$ generated in Equation (1), a dedicated Code Generation LLM produces the corresponding visualization code using chain-of-thought (CoT) prompting:
$C_{\_}i=f_{\_}{\text{code}}(D,R_{\_}i),$ (2)
where $f_{\_}{\text{code}}$ represents the code generation function. Unlike naive code generation approaches, here, the dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes. The generated code $C_{\_}i$ is then executed to render a candidate visualization:
$V_{\_}i=f_{\_}{\text{exec}}(C_{\_}i),\quad i=1,2,\dots,K,$ (3)
with $f_{\_}{\text{exec}}$ serving as the code execution function. By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage. In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
$\epsilon_{\_}i=\begin{cases}1,&\text{if }C_{\_}i\text{ is executable}\\ 0,&\text{otherwise}.\end{cases}$ (4)
To route the outputs appropriately, we introduce:
$Z_{\_}i=\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1,\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0.\end{cases}$ (5)
The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the original query $Q$, to the feedback model in the subsequent stage.
3.3 Feedback-Driven Code Optimization
Most code generation frameworks focus merely on syntactically correct scripts, but our framework goes further. As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback. A Vision-Language Model (VLM) is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$. This evaluation is formalized as:
$F_{\_}i=f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i),$ (6)
where $F_{\_}i$ provides structured feedback on key aspects such as chart layout, the alignment between the intended request and the rendered visualization (or error context), and visual readability (including potential improvements). To capture the complete quality signal from each candidate, we pair the feedback with its corresponding generated code:
$S_{\_}i=(C_{\_}i,F_{\_}i),\quad i=1,2,\dots,K.$ (7)
Leveraging the collective code-feedback pairs along with the original query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
$C^{*}=f_{\_}{\text{integrate}}\left(Q,D,\{S_{\_}i\}_{\_}{i=1}^{K}\right),$ (8)
where $C^{*}$ represents the optimized visualization code and $f_{\_}{\text{integrate}}$ denotes the function that aggregates the strengths of each candidate code alongside its corresponding feedback. This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description. The process harnesses the reasoning capabilities of LLM to systematically evaluate the strengths and pinpoint potential weaknesses across all candidate solutions, including those that may initially be non-executable. By synthesizing the most optimal elements from each candidate, the framework generates a final code that best captures the user’s intent while maintaining high standards of reliability, robustness, and execution quality. A step-by-step breakdown of this process is detailed in Algorithm 1 below:
Algorithm 1 Algorithm for VisPath
1:User query $Q$, dataset $D$, number of reasoning paths $K$, Multi-Path Agent $f_{\_}{\text{mpa}}$, Code Generation LLM $f_{\_}{\text{code}}$, Code Execution Function $f_{\_}{\text{exec}}$, Feedback Model $f_{\_}{\text{feedback}}$, Integration Module $f_{\_}{\text{integrate}}$
2:Final visualization code $C^{*}$
3:// Step 1: Multi-Path Query Expansion
4:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$triangleright Generate $K$ distinct reasoning paths; each $R_{\_}i$ outlines a potential interpretation of $Q$ given $D$.
5:// Step 2: Code Generation from Reasoning Paths
6:for $i=1$ to $K$ do
7:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$triangleright Generate candidate visualization code $C_{\_}i$ based on reasoning path $R_{\_}i$.
8:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$triangleright Execute $C_{\_}i$ to obtain candidate visualization $V_{\_}i$.
9:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$triangleright Record executability indicator $\epsilon_{\_}i$ for $C_{\_}i$.
10:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$triangleright Route output: either the rendered visualization or the error message.
11:end for
12:// Step 3: Feedback-Driven Code Optimization
13:for $i=1$ to $K$ do
14:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$triangleright Obtain structured feedback $F_{\_}i$ on candidate code $C_{\_}i$ using the routed output $Z_{\_}i$.
15:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$triangleright Form a unified tuple $S_{\_}i$ combining $C_{\_}i$, $F_{\_}i$.
16:end for
17:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$triangleright Aggregate the code-feedback pairs along with $Q$ and $D$ to synthesize the final visualization code $C^{*}$.
18:return $C^{*}$
4 Experiments
Model Methods MatPlotBench Qwen-Agent Code Interpreter benchmark
Plot Score Executable Rate (%) Visualization-Hard Visualization-Easy Avg.
GPT-4o mini Zero-Shot 62.38 53 59.68 45.50 52.59
CoT Prompting Wei et al. (2022) 61.95 50 57.50 40.00 48.75
Chat2VIS Maddigan and Susnjak (2023) 56.98 53 59.36 36.50 47.93
MatPlotAgent Yang et al. (2024) 63.90 58 67.50 53.25 60.38
VisPath (Ours) 66.12 60 70.68 57.23 63.96
Gemini 2.0 Flash Zero-Shot 55.00 54 68.97 52.18 60.58
CoT Prompting Wei et al. (2022) 53.56 61 40.00 63.89 51.95
Chat2VIS Maddigan and Susnjak (2023) 54.89 55 59.36 56.50 57.93
MatPlotAgent Yang et al. (2024) 56.31 58 77.62 51.50 64.56
VisPath (Ours) 59.37 63 80.79 57.17 68.98
Table 1: Performance comparison of various methods across different benchmarks. Zero-Shot refers to directly generating code. CoT Prompting utilizes Chain of Thought Prompting. Visualization-Hard and Visualization-Easy refer to the Accuracy of Code Execution Results on different subsets of the Qwen-Agent Code Interpreter benchmark. Bold text indicates the best performance, underlined text indicates the second-best performance. denotes our proposed method.
4.1 Setup
In this section, we detail our experimental configuration, including experimental datasets, model specifications, and baseline methods for evaluating the performance of the proposed VisPath framework in generating visualization code from natural language queries.
4.1.1 Experimental Datasets
We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al. (2024) and the Qwen-Agent Code Interpreter Benchmark. MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries. In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
4.1.2 Models Used
Large Language Models (LLMs):
For the code inference stage, we experiment with GPT-4o mini Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) to generate candidate visualization code from the reasoning paths. Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with recommendations from previous work Yang et al. (2024). To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al. (2023) and Gemini 2.0 Flash Team et al. (2024) as our visualization feedback model, which provides high-quality reference assessments.
Vision Language Models (VLMs):
In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework. Specifically, GPT-4o Achiam et al. (2023) is employed for detailed plot evaluation in all evaluation tasks. This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
4.1.3 Evaluation Metrics
In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability. MatPlotBench assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions. Qwen-Agent Code Interpreter benchmark evaluates visualization models based on Visualization-Hard and Visualization-Easy, measuring how well generated images align with queries of different difficulty levels. Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric. Previous studies show GPT-based VLM evaluations align with human assessments, hence VLM was used for evaluation.
4.1.4 Baseline Methods
We compare VisPath against competitive baselines. Specifically, (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses chain-of-thought prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al. (2024) first expands the query and then refines the code via a self-debugging loop with feedback. For a fair comparison, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well, and VisPath generates three reasoning paths with corresponding visual feedback to refine the final output. [^2] Prompts are detailed in Appendix A.
4.2 Experimental Analysis
In this section, we present a detailed evaluation of our proposed VisPath framework, summarizing the performance results against four baselines: (1) Zero-Shot, (2) Chain-of-Thought (CoT) Prompting, (3) Chat2VIS, and (4) MatPlotAgent.
From Table  1, we find that the Zero-Shot method, which directly generates visualization code from the user query without intermediate reasoning, suffers from ambiguous interpretations and error-prone outputs. These issues are particularly pronounced when the input queries are underspecified, contain implicit assumptions, or require complex reasoning steps. CoT Prompting, on the other hand, mitigates some of these challenges by employing chain-of-thought (CoT) reasoning, which helps the model articulate its decision-making process step by step. While this structured reasoning approach improves interpretability and correctness, its reliance on a single reasoning trajectory limits its capacity to explore alternative solutions.
Meanwhile, Chat2VIS builds upon CoT Prompting by incorporating guided prefixes to better structure user queries and disambiguate input intent. This enhancement leads to more coherent code generation and reduces errors stemming from unclear specifications. However, its effectiveness still depends on the predefined guiding templates, which may not fully adapt to highly variable or novel query structures. MatPlotAgent further refines this approach through query expansion and an iterative self-debugging loop that enhances code robustness. However, its reliance on an iterative correction mechanism comes at the cost of computational efficiency and does not fully leverage diverse reasoning strategies that could lead to more creative and contextually accurate solutions.
In contrast, our novel framework, VisPath, overcomes these challenges by dynamically generating multiple reasoning paths. By exploring diverse interpretations of user intent simultaneously, VisPath significantly enhances the robustness of the generated visualization code. Furthermore, it integrates structured feedback from Vision-Language Models (VLMs) to refine its outputs, ensuring both higher execution accuracy and improved aesthetic quality. This feedback-driven optimization resolves ambiguities more effectively and selects the most appropriate solution from a wide range of possibilities. Evaluations conducted on benchmark datasets such as MatPlotBench and Qwen-Agent Code Interpreter benchmark demonstrate that VisPath consistently outperforms the baseline methods by average 17%. VisPath exhibits a superior ability to handle complex and ambiguous visualization requests by synthesizing multiple reasoning trajectories and incorporating structured feedback, making it a highly effective tool for automated visualization generation.
4.3 Ablation Study
This section further explores the impact of varying the number of reasoning paths on the performance of visualization code generation. We conduct a series of ablation studies to demonstrate the robustness of our VisPath framework under alternative settings. Specifically, we investigate (i) the effect of varying the number of generated reasoning paths and (ii) the impact of a simplified integration strategy for synthesizing the final visualization code.
4.3.1 Varying the Number of Reasoning Paths
To further assess the contribution of reasoning path diversity, we conducted ablation experiments varying the number of generated paths $K\in\{2,3,4\}$ to evaluate their effects on execution accuracy, visualization quality and interpretability. Evaluation has been conducted on both the MatPlotBench and Qwen-Agent Code Interpreter benchmark datasets. The study reveals an interesting pattern observed when the number of reasoning path is reduced. As illustrated in Table 2 and Table 3, when $K=2$, the system exhibited limited diversity in which occasionally results in missing nuanced interpretations of complex user queries.
Model K MatPlotBench
Plot Score Executable Rate (%)
GPT-4o mini 2 64.02 58
3 66.12 60
4 64.68 62
Gemini 2.0 Flash 2 56.59 56
3 59.37 61
4 54.60 54
Table 2: Results of ablation study on MatPlotBench.
Model K Qwen-Agent
Hard Easy Avg.
GPT-4o mini 2 67.50 55.72 61.61
3 70.68 57.23 63.96
4 66.42 50.25 58.34
Gemini 2.0 Flash 2 74.21 57.50 65.86
3 80.79 57.17 68.98
4 76.41 54.50 65.46
Table 3: Results of ablation study on Qwen-Agent Code Interpreter benchmark.
Conversely, increasing $K$ to 4 introduces extra paths which can sometimes add noise as less relevant interpretations are considered. Specifically, we observed cases where redundant or overly complex visualization components were generated, hence making the final output harder to be interpreted. While the additional reasoning paths increased the overall exploration space, they also required more extensive filtering and selection, leading to suboptimal execution efficiency.
Notably, $K=3$ consistently achieves the best balance by providing sufficient diversity without introducing excessive noise. With too few reasoning paths, some nuanced aspects of ambiguous queries are lost, while too many paths configuration result in unnecessary exploration of low relevancy solutions. The number of reasoning path diversity set as $K=3$ optimally provides sufficient diversity without overcomplicating the reasoning process.
4.3.2 Robustness with a Simple Integration Strategy
To further validate the robustness of VisPath, we also evaluate an alternative integration strategy that simplicies the aggregation of multiple reasoning paths. In contrast to our full feedback-driven iterative approach which refines visualization code through multiple rounds of vision-language feedback, this streamlined method directly aggregates candidate outputs from different reasoning trajectories without intermediate corrections.
Specifically, in this alternative strategy, three candidate codes, each derived from different reasoning paths, are directly combined to generate the final visualization output. This methodology reduces computational overhead and processing time while maintaining a degree of interpretability and correctness. Through examining the impact of this approach, we gain insight into the extent to which VisPath’s core strength stems from its multi-path reasoning capability versus its iterative refinement process.
Model Feedback MatPlotBench Qwen-Agent
Plot Score Executable Rate (%) Avg.
GPT-4o mini (w/o) feedback 63.76 56 58.00
(w) feedback 66.12 60 63.96
Gemini 2.0 Flash (w/o) feedback 55.28 57 64.03
(w) feedback 59.37 63 68.98
Table 4: Performance comparison of VisPath with and without visual feedback. The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two llm models.
As shown in Table 4, even under this simplified integration, the performance of VisPath significantly outperforms all baseline methods, demonstrating its robustness and adaptability. This result suggests that the diversity of reasoning paths alone ensures a resilient and effective outcome, even in the absence of extensive intermediate corrections. This finding shows that the strength of our framework largely stems from its multi-path reasoning design. Even when the integration is simplified, the diversity of reasoning paths ensures that the aggregated output remains robust and effective, thus showing the overall adaptability and effectiveness of our approach.
5 Discussion
Our main experiment and ablation studies decisively demonstrate that VisPath’s multi-path exploration revolutionizes visualization code generation by making it not just robust, but truly interpretable.  [^3] Case study is provided in Appendix B. Unlike traditional methods that often yield simplistic, one-dimensional results, VisPath embraces complexity—delving deeper to capture nuances of user intent that would otherwise be lost. By integrating multiple perspectives during reasoning, our framework ensures that final visualizations are not only data-accurate but also rich in context, inherently intuitive, and unmistakably aligned with the user’s true objectives.
The real breakthrough with VisPath is how it transforms visualization into a structured yet fluid process. Traditional methods often act like static translators, converting user queries into plots while neglecting the vital auxiliary components—legends, labels, line styles, and other elements—that enable true interpretability. They assume that a single pass can sufficiently capture the essence of a visualization, but this leads to visualizations that might technically be correct but fail at communication. In contrast, VisPath’s multi-path approach systematically explores different ways to represent data, integrating feedback at each step. This ensures that even subtle yet critical design elements are not just included, but optimized, making the resulting visualizations feel complete, polished, and immediately comprehensible.
Beyond accuracy, VisPath introduces a level of adaptability that is indispensable in today’s fast-evolving data landscape. It’s not just a tool for static reports—it is built for real-time streaming data, dynamic dashboards, and interactive analytics. Even with broad or loosely defined prompts, our framework intelligently synthesizes high-quality visualizations, closing the gap between intent and execution. As industries continue to demand clearer, faster, and more adaptable insights, VisPath is poised to lead the way, setting a new standard for intelligent, context-aware data visualization.
6 Conclusion
In this work, we present VisPath, a groundbreaking framework that redefines the landscape of automated visualization code generation over existing methods. Unlike prior methods, our approach seaminglessly combines multi-path reasoning with feedback-driven optimization. Accurately capturing diverse user intents and refine generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark. By prioritizing adapability and user intent alignment, VisPath is uniquely positioned to handle compledities of real-world data visualization tasks. Looking ahead, future work could explore VisPath’s adaptability in more dynamic, real-world visualization scenarios, further broadening its scope and enhancing its practical utility in complex data analysis contexts.
7 Limitations
Despite its effectiveness, the current iteration of our framework relies on a limited feedback mechanism, evaluating visualizations primarily based on two aspects: query-code alignment and query-plot image alignment. While these aspects provide valuable insights, they may not fully capture the granular elements that contribute to overall interpretability. To enhance the depth of feedback itself, future research could focus on decomposition of assessment process, assessing individual plot components separately. This granular analysis would render for a more nuanced assessment of readability, element appropriateness and visual coherence, ultimately leading to more refined visualization code generation. Strengthening the feedback mechanism in such way will be crucial for maximizing VisPath’s effectiveness in diverse and complex visualization scenarios.
Appendix A Appendix A. Prompts Used
Prompt for Multi-Path Reasoning [System Prompt] According to the user query, expand and solidify the query into detailed instruction on how to write python code to fulfill the user query’s requirements. Import the appropriate libraries. Pinpoint the correct library functions to call and set each parameter in every function call accordingly. [User Prompt] Think step by step. Generate three distinct extended queries based on the given query. Ensure that you first analyze the given data description and create queries that align with the data. If no data description is provided, follow the original query as is. You must follow the Python list format for the output. Do not modify the detailed instructions from the original user query. Original query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
Prompt for Code Generation [System Prompt] You are an expert in data visualization code generation. Think step by step and write the generated code in the format “‘python…“‘, where "…" represents the generated code. The code must end with ‘plt.show()‘. [User Prompt] Think step by step. Based on the user’s query and the provided data description, generate Python code using ‘matplotlib.pyplot‘ and ‘seaborn‘ to create the requested plot. Ensure that the code is formatted as “‘…“‘, where "…" represents the generated code. User query: {query} Data description: {datadescription}
Prompt for Visual Feedback [System Prompt] Given a code, a user query, and an image of the current plot, please determine whether the plot accurately follows the user query. Provide detailed instructions on how to enhance the plot using Python code. [User Prompt] Carefully analyze the provided Python code, the user query, and the plot image (if available) to assess whether the generated plot meets the user query requirements. If the plot image is missing, check the error message that occurred in the code. Compare the plot with the user query, highlight discrepancies, and provide clear, actionable instructions to modify the code. Additionally, suggest improvements for better visualization, focusing on clarity, readability, and alignment with the user’s objectives. Code: {code} User query: {oriquery}
Prompt for Synthesis [System Prompt] You are an expert on data visualization code judgement and aggregation. [User Prompt] Think step by step. Given the provided user query, data description, multiple data visualization codes generated for the query, and feedback for each code’s generated image. Your task is to: 1. Carefully review the user query and the data description. 2. Examine each version of the data visualization code along with the feedback provided for each version. 3. Synthesize the feedbacks for each code, user query insights, data description to create a final version of the code. 4. Your goal is to produce a final version of code that more effectively fulfills the user query by integrating the best elements from all versions and applying necessary corrections. User Query: {oriquery} Data Description: {datadescription} Code for aggregation with corresponding feedback: {codeforaggregation}
Prompt for Evaluation: MatplotBench You are an excellent judge at evaluating visualization plots between a model generated plot and the ground truth. You will be giving scores on how well it matches the ground truth plot. The generated plot will be given to you as the first figure. If the first figure is blank, that means the code failed to generate a figure. Another plot will be given to you as the second figure, which is the desired outcome of the user query, meaning it is the ground truth for you to reference. Please compare the two figures head to head and rate them. Suppose the second figure has a score of 100, rate the first figure on a scale from 0 to 100. Scoring should be carried out in the following aspect: 1. Plot correctness: Compare closely between the generated plot and the ground truth, the more resemblance the generated plot has compared to the ground truth, the higher the score. The score should be proportionate to the resemblance between the two plots. In some rare occurrence, see if the data points are generated randomly according to the query, if so, the generated plot may not perfectly match the ground truth, but it is correct nonetheless. Only rate the first figure, the second figure is only for reference. If the first figure is blank, that means the code failed to generate a figure. Give a score of 0 on the Plot correctness. After scoring from the above aspect, please give a final score. The final score is preceded by the [FINAL SCORE] token. For example [FINAL SCORE]: 40.
Prompt for Evaluation: Qwen-Agent Code Interpreter benchmark Please judge whether the image is consistent with the [Question] below, if it is consistent then reply "right", if not then reply "wrong". Question: {query}
Appendix B Appendix B. Case Study
Figure 3: Overview of end-to-end process of visualizing women’s million-dollar earnings data using VisPath. It includes the user query, raw data, multi-path reasoning, generated figures, visual feedback, final script, and the completed visualization.

### v2

@@
-VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
+Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
 Abstract
-Unprecedented breakthroughs in Large Language Models (LLMs) has amplified its penetration into application of automated visualization code generation.
-Few-shot prompting and query expansion techniques have notably enhanced data visualization performance, however, still fail to overcome ambiguity and complexity of natural language queries - imposing an inherent burden for manual human intervention.
-To mitigate such limitations, we propose a holistic framework VisPath : A Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation, which systematically enhances code quality through structured reasoning and refinement.
-VisPath is a multi-stage framework, specially designed to handle underspecified queries.
-To generate a robust final visualization code, it first utilizes initial query to generate diverse reformulated queries via Chain-of-Thought (CoT) prompting, each representing a distinct reasoning path.
-Refined queries are used to produce candidate visualization scripts, consequently executed to generate multiple images.
-Comprehensively assessing correctness and quality of outputs, VisPath generates feedback for each image, which are then fed to aggregation module to generate optimal result.
-Extensive experiments on benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath significantly outperforms state-of-the-art (SOTA) methods, increased up to average 17%, offering a more reliable solution for AI-driven visualization code generation.
-VisPath: Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
-Wonduk Seo1thanks: Equal contribution.
-Seungyong Lee2*   Daye Kang1,3   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author.
-1Enhans.ai   2Unaffiliated   3KAIST   4Tsinghua University 1{seowonduk,seunghyun,daye}@enhans.ai 2leesy7197@khu.ac.kr   4yzh23@tsinghua.edu.cn
+Rapid advancements in Large Language Models (LLMs) have accelerated their integration into automated visualization code generation applications.
+Despite advancements through few-shot prompting and query expansion, existing methods remain limited in handling ambiguous and complex queries, thereby requiring manual intervention.
+To overcome these limitations, we propose VisPath: a Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation.
+VisPath handles underspecified queries through structured, multi-stage processing.
+It begins by reformulating the user input via Chain-of-Thought (CoT) prompting, which refers to the initial query while generating multiple extended queries in parallel, enabling the LLM to capture diverse interpretations of the user intent.
+These queries then generate candidate visualization scripts, which are executed to produce diverse images.
+By assessing the visual quality and correctness of each output, VisPath generates targeted feedback that is aggregated to synthesize an optimal final result.
+Extensive experiments on widely-used benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, offering a more reliable solution for AI-driven visualization code generation.
+Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
+Wonduk Seo1,2   Seungyong Lee   Daye Kang1,3 Hyunjin An1   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author.
+1Enhans   2Peking University   3KAIST   4Tsinghua University 1{wonduk,daye,hyunjin,seunghyun}@enhans.ai
 1 Introduction
@@
 (2005); Bisong and Bisong (2019); Zhu (2013).
-This approach demands programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al.
+These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al.
 (2018); Sharif et al.
@@
 Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation.
-However, these methods face several limitations: ① both generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; ② both rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries.
-③ both approaches encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback.
-Without a mechanism to retrieve outputs that reflect diverse possibilities, they struggle to capture intricate details, ultimately limiting the precision and adaptability of the generated visualizations.
-To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a transformative approach that redefines how visualization code is generated.
-Traditional single-pass methods may seem efficient, but they often fall short of delivering the depth and precision users truly need.
-They generate code quickly, but struggle to capture the intricate details that make a visualization not just functional but meaningful.
+However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; (2) they rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries; (3) they encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback.
+Without a mechanism to retrieve outputs that reflect diverse possibilities, they face difficulties in capturing the intricate details required for visualizations that are both functionally precise and contextually relevant.
+To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a novel approach that redefines how visualization code is generated, as illustrated in Figure 1.
+Traditional methods often fall short of delivering the depth and precision users truly need, struggle to capture the intricate details that make a visualization not just functional but meaningful.
 VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
-Rather than simply translating user input into code, VisPath ensures that every critical aspect—both explicitly stated and implicitly necessary—is carefully considered, creating a visualization that is not only correct but insightful.
+Rather than simply translating user input into code, VisPath ensures that every critical aspect, both explicitly stated and implicitly necessary, is carefully considered, creating a visualization that is not only correct but insightful.
 At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting.
@@
 The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact.
-By shifting from a single-pass, one-size-fits-all approach to a dynamic, multi-layered process, VisPath sets a new standard for visualization generation—one that is not just about creating code, but about ensuring data is represented in its most clear, meaningful, and compelling form.
 Extensive experiments on benchmark datasets, including MatPlotBench Yang et al.
-(2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate VisPath’s superiority over the state-of-the-art (SOTA) visualization generation methods.
-By systematically generating and evaluating multiple reasoning paths and leveraging iterative feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents.
+(2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate that VisPath notably outperforms state-of-the-art methods across all evaluation metrics.
+In addition, ablation studies confirm that VisPath’s performance gains primarily stem from its Multi-Path Reasoning mechanism and feedback-driven optimization.
+By systematically generating and evaluating multiple reasoning paths and leveraging feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents.
 Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
@@
 The framework consists of combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
-Building on these machine learning strategies, numerous studies have utilized Large Language Models (LLMs) to further enhance system flexibility.
+Building on these machine learning strategies, numerous studies have utilized LLMs to further enhance system flexibility.
 Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al.
@@
 By committing to a single reasoning trajectory, they may produce code that is syntactically correct yet semantically misaligned with user expectations, requiring extensive manual adjustments.
-This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification—ironically undermining the intended efficiency of automation.
-To address these limitations, we introduce VisPath, a novel framework that integrates multi-path reasoning with feedback from Vision-Language Models (VLMs) to enhance visualization code generation.
-Compared to conventional methods constrained by a single reasoning path, VisPath dynamically explores multiple solution pathways, improving interpretability, minimizing manual intervention, and adapting more effectively to diverse visualization requirements.
+This challenge is particularly pronounced when user input is ambiguous or underspecified, leading to an iterative cycle of prompt refinement and code modification, consequently limiting the intended efficiency of automation.
+To address these limitations, we introduce VisPath, a novel framework that integrates Multi-Path Reasoning with feedback from VLMs to enhance visualization code generation.
 3 Methodology
 We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback.
-VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a Vision-Language Model (VLM) evaluates and refines the outputs to ensure generation robust visualization code.
+VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a VLM evaluates and refines the outputs to ensure generation robust visualization code.
 An overview of this process is shown in Figure 2.
 3.1 Multi-Path Generation
-One of the biggest pitfalls in visualization code generation is rigid interpretation, a single query can have multiple valid visual representations depending on its dataset structure.
+The potential for rigid interpretation is a major limitation in visualization code generation, a single query can have multiple valid representations depending on its dataset structure.
 VisPath mitigates this by generating multiple extended queries within a single interaction.
@@
 Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request.
-This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent—even when queries are ambiguous or underspecified.
+The purpose of generating multiple reasoning paths is not simply to increase their quantity, but to ensure interpretive diversity.
+Vispath mitigates the risk of depending on a single, potentially incorrect assumption by intentionally exploring alternative and context-aware reasoning strategies.
+This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent even when queries are ambiguous or underspecified.
 3.2 Code Generation from Reasoning Paths
-Once diverse reasoning paths are established, the next stage involves translating each path into executable Python scripts.
-For every reasoning path $R_{\_}i$ generated in Equation (1), a dedicated Code Generation LLM produces the corresponding visualization code using chain-of-thought (CoT) prompting:
+Once diverse reasoning paths are established, the subsequent stage involves translating each path into executable Python scripts.
+For each reasoning path $R_{\_}i$ generated in Equation 1, a dedicated Code Generation LLM produces the corresponding visualization code using Chain-of-Thought (CoT) prompting:
 $C_{\_}i=f_{\_}{\text{code}}(D,R_{\_}i),$ (2)
 where $f_{\_}{\text{code}}$ represents the code generation function.
-Unlike naive code generation approaches, here, the dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes.
+The dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes.
 The generated code $C_{\_}i$ is then executed to render a candidate visualization:
@@
 with $f_{\_}{\text{exec}}$ serving as the code execution function.
-By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage.
-In practice, some generated codes may not be executable.
-Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
+By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage, as the code has already been conditioned on $D$ during its generation.
+In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
 $\epsilon_{\_}i=\begin{cases}1,&\text{if }C_{\_}i\text{ is executable}\\ 0,&\text{otherwise}.\end{cases}$ (4)
@@
 $Z_{\_}i=\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1,\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0.\end{cases}$ (5)
-The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the original query $Q$, to the feedback model in the subsequent stage.
+The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the initial query $Q$, to the feedback model in the subsequent stage.
 3.3 Feedback-Driven Code Optimization
-Most code generation frameworks focus merely on syntactically correct scripts, but our framework goes further.
+While most code generation frameworks primarily focus on producing syntactically correct scripts, our framework extends beyond this by incorporating an additional mechanism.
 As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback.
-A Vision-Language Model (VLM) is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$.
+A VLM is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$.
 This evaluation is formalized as:
@@
 $S_{\_}i=(C_{\_}i,F_{\_}i),\quad i=1,2,\dots,K.$ (7)
-Leveraging the collective code-feedback pairs along with the original query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
+Leveraging the collective code-feedback pairs along with the initial query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
 $C^{*}=f_{\_}{\text{integrate}}\left(Q,D,\{S_{\_}i\}_{\_}{i=1}^{K}\right),$ (8)
@@
 This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description.
-The process harnesses the reasoning capabilities of LLM to systematically evaluate the strengths and pinpoint potential weaknesses across all candidate solutions, including those that may initially be non-executable.
-By synthesizing the most optimal elements from each candidate, the framework generates a final code that best captures the user’s intent while maintaining high standards of reliability, robustness, and execution quality.
-A step-by-step breakdown of this process is detailed in Algorithm 1 below:
+An algorithm for VisPath is detailed in Algorithm 1 as shown below:
 Algorithm 1 Algorithm for VisPath
 1:User query $Q$, dataset $D$, number of reasoning paths $K$, Multi-Path Agent $f_{\_}{\text{mpa}}$, Code Generation LLM $f_{\_}{\text{code}}$, Code Execution Function $f_{\_}{\text{exec}}$, Feedback Model $f_{\_}{\text{feedback}}$, Integration Module $f_{\_}{\text{integrate}}$
-2:Final visualization code $C^{*}$
-3:// Step 1: Multi-Path Query Expansion
-4:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$triangleright Generate $K$ distinct reasoning paths; each $R_{\_}i$ outlines a potential interpretation of $Q$ given $D$.
-5:// Step 2: Code Generation from Reasoning Paths
-6:for $i=1$ to $K$ do
-7:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$triangleright Generate candidate visualization code $C_{\_}i$ based on reasoning path $R_{\_}i$.
-8:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$triangleright Execute $C_{\_}i$ to obtain candidate visualization $V_{\_}i$.
-9:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$triangleright Record executability indicator $\epsilon_{\_}i$ for $C_{\_}i$.
-10:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$triangleright Route output: either the rendered visualization or the error message.
-11:end for
-12:// Step 3: Feedback-Driven Code Optimization
-13:for $i=1$ to $K$ do
-14:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$triangleright Obtain structured feedback $F_{\_}i$ on candidate code $C_{\_}i$ using the routed output $Z_{\_}i$.
-15:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$triangleright Form a unified tuple $S_{\_}i$ combining $C_{\_}i$, $F_{\_}i$.
-16:end for
-17:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$triangleright Aggregate the code-feedback pairs along with $Q$ and $D$ to synthesize the final visualization code $C^{*}$.
-18:return $C^{*}$
+2:// Step 1: Multi-Path Query Expansion
+3:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$
+4:// Step 2: Code Generation from Reasoning Paths
+5:for $i=1$ to $K$ do
+6:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$
+7:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$
+8:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$
+9:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$
+10:end for
+11:// Step 3: Feedback-Driven Code Optimization
+12:for $i=1$ to $K$ do
+13:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$
+14:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$
+15:end for
+16:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$
+17:return $C^{*}$
 4 Experiments
+4.1 Setup
+In this section, we detail our experimental configuration, including (1) experimental datasets, (2) model specifications, and (3) baseline methods for evaluating the performance of the proposed VisPath.
+4.1.1 Experimental Datasets
+We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al.
+(2024) and Qwen-Agent Code Interpreter Benchmark.
+Specifically, MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries.
+In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
+4.1.2 Models Used
+Large Language Models (LLMs):
+For the code inference stage, we experiment with GPT-4o mini Achiam et al.
+(2023) and Gemini 2.0 Flash Team et al.
+(2024) to generate candidate visualization code from the reasoning paths.
+Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work Yang et al.
+(2024).
+To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al.
+(2023) and Gemini 2.0 Flash Team et al.
+(2024) as our visualization feedback model, which provides high-quality reference assessments.
+Vision-Language Models (VLMs):
+In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework.
+Specifically, GPT-4o Achiam et al.
+(2023) is employed for detailed plot evaluation in all evaluation tasks.
+This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
+4.1.3 Evaluation Metrics
+In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability.
+MatPlotBench Bisong and Bisong (2019) assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions.
+Qwen-Agent Code Interpreter benchmark [^2] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy.
+Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric.
+Previous studies showed that GPT-based VLM evaluations align well with human assessments Yang et al.
+(2024), hence VLM was used for evaluation.
 Model Methods MatPlotBench Qwen-Agent Code Interpreter benchmark
@@
 denotes our proposed method.
-4.1 Setup
-In this section, we detail our experimental configuration, including experimental datasets, model specifications, and baseline methods for evaluating the performance of the proposed VisPath framework in generating visualization code from natural language queries.
-4.1.1 Experimental Datasets
-We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al.
-(2024) and the Qwen-Agent Code Interpreter Benchmark.
-MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries.
-In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
-4.1.2 Models Used
-Large Language Models (LLMs):
-For the code inference stage, we experiment with GPT-4o mini Achiam et al.
-(2023) and Gemini 2.0 Flash Team et al.
-(2024) to generate candidate visualization code from the reasoning paths.
-Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with recommendations from previous work Yang et al.
-(2024).
-To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al.
-(2023) and Gemini 2.0 Flash Team et al.
-(2024) as our visualization feedback model, which provides high-quality reference assessments.
-Vision Language Models (VLMs):
-In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework.
-Specifically, GPT-4o Achiam et al.
-(2023) is employed for detailed plot evaluation in all evaluation tasks.
-This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
-4.1.3 Evaluation Metrics
-In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability.
-MatPlotBench assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions.
-Qwen-Agent Code Interpreter benchmark evaluates visualization models based on Visualization-Hard and Visualization-Easy, measuring how well generated images align with queries of different difficulty levels.
-Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric.
-Previous studies show GPT-based VLM evaluations align with human assessments, hence VLM was used for evaluation.
 4.1.4 Baseline Methods
-We compare VisPath against competitive baselines.
-Specifically, (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses chain-of-thought prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al.
+We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al.
 (2024) first expands the query and then refines the code via a self-debugging loop with feedback.
-For a fair comparison, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well, and VisPath generates three reasoning paths with corresponding visual feedback to refine the final output.
+Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output.
 [^3] Prompts are detailed in Appendix A.
+For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well.
 4.2 Experimental Analysis
-In this section, we present a detailed evaluation of our proposed VisPath framework, summarizing the performance results against four baselines: (1) Zero-Shot, (2) Chain-of-Thought (CoT) Prompting, (3) Chat2VIS, and (4) MatPlotAgent.
-From Table  1, we find that the Zero-Shot method, which directly generates visualization code from the user query without intermediate reasoning, suffers from ambiguous interpretations and error-prone outputs.
-These issues are particularly pronounced when the input queries are underspecified, contain implicit assumptions, or require complex reasoning steps.
-CoT Prompting, on the other hand, mitigates some of these challenges by employing chain-of-thought (CoT) reasoning, which helps the model articulate its decision-making process step by step.
-While this structured reasoning approach improves interpretability and correctness, its reliance on a single reasoning trajectory limits its capacity to explore alternative solutions.
-Meanwhile, Chat2VIS builds upon CoT Prompting by incorporating guided prefixes to better structure user queries and disambiguate input intent.
-This enhancement leads to more coherent code generation and reduces errors stemming from unclear specifications.
-However, its effectiveness still depends on the predefined guiding templates, which may not fully adapt to highly variable or novel query structures.
-MatPlotAgent further refines this approach through query expansion and an iterative self-debugging loop that enhances code robustness.
-However, its reliance on an iterative correction mechanism comes at the cost of computational efficiency and does not fully leverage diverse reasoning strategies that could lead to more creative and contextually accurate solutions.
-In contrast, our novel framework, VisPath, overcomes these challenges by dynamically generating multiple reasoning paths.
-By exploring diverse interpretations of user intent simultaneously, VisPath significantly enhances the robustness of the generated visualization code.
-Furthermore, it integrates structured feedback from Vision-Language Models (VLMs) to refine its outputs, ensuring both higher execution accuracy and improved aesthetic quality.
-This feedback-driven optimization resolves ambiguities more effectively and selects the most appropriate solution from a wide range of possibilities.
-Evaluations conducted on benchmark datasets such as MatPlotBench and Qwen-Agent Code Interpreter benchmark demonstrate that VisPath consistently outperforms the baseline methods by average 17%.
-VisPath exhibits a superior ability to handle complex and ambiguous visualization requests by synthesizing multiple reasoning trajectories and incorporating structured feedback, making it a highly effective tool for automated visualization generation.
+VisPath is evaluated against four baselines: Zero Shot prompting, CoT prompting, Chat2VIS, and MatPlotAgent.
+The evaluation is conducted on MatPlotBench and the Qwen-Agent Code Interpreter benchmark using GPT-4o mini and Gemini 2.0 Flash, as shown in Table 1.
+Zero-Shot prompting generates visualization code directly from natural language queries without intermediate reasoning.
+While computationally efficient, it often struggles to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs.
+On MatPlotBench (GPT-4o mini), it achieves a Plot Score of 62.38 and an Executable Rate of 53%.
+CoT prompting further introduces a single reasoning step to expose intermediate decisions and improve interpretability.
+However, on MatPlotBench, it slightly underperforms Zero-Shot in both Plot Score and Executable Rate, indicating reliance on a fixed reasoning path may reduce adaptability to diverse input structures.
+Chat2VIS extends CoT prompting by adopting prefix templates to improve coherence and reduce ambiguity in user instructions.
+While this approach is effective for well-structured or common query formats, its dependence on fixed templates limits adaptability when processing loosely specified or novel queries.
+Such limitation is evident in its performance on MatPlotBench, where it achieves a Plot Score of 56.98 and an Executable Rate of 53%.
+Furthermore, MatPlotAgent incorporates query expansion and iterative self-debugging mechanisms to enhance robustness.
+While effective at correcting execution-level errors, its revisions are confined to localized adjustments and do not address higher-order semantic ambiguities.
+In contrast, our proposed framework VisPath is specifically designed to overcome these limitations observed in prior methods by dynamically generating multiple reasoning paths and refining them through structured visual feedback.
+In particular, template-based approaches such as Chat2VIS offer limited adaptability due to their reliance on predefined input formats, while methods such as MatPlotAgent focus on localized corrections without addressing broader semantic ambiguity.
+Unlike prior methods, VisPath generates diverse interpretations of user intent and evaluates them holistically using structured vision-language feedback.
+This enables more flexible handling of under-specified or ambiguous inputs, resulting in semantically aligned and executable visualizations.
+Evaluated across multiple benchmark settings, VisPath notably outperforms baselines, achieving up to 9.14 point gains in Plot Score and a 10% point increase in Executable Rate.
+These improvements well demonstrate VisPath’s robustness in exploring diverse reasoning paths and iteratively refine outputs through structured visual feedback, effectively reducing semantic ambiguity and improving execution reliability.
 4.3 Ablation Study
-This section further explores the impact of varying the number of reasoning paths on the performance of visualization code generation.
-We conduct a series of ablation studies to demonstrate the robustness of our VisPath framework under alternative settings.
-Specifically, we investigate (i) the effect of varying the number of generated reasoning paths and (ii) the impact of a simplified integration strategy for synthesizing the final visualization code.
+To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments in this section.
+Specifically, we analyze the following three aspects: (i) varying the number of generated reasoning paths, (ii) the effect of removing visual feedback during integration, and (iii) the contribution of visual feedback beyond binary executability.
+Figure 3: Effect of varying the number of reasoning paths $K$ on performance across datasets and models.
+Metrics include Plot Score, Executable Rate.
+The results show that $K=3$ achieves the best overall balance, with larger $K$ values reducing performance.
 4.3.1 Varying the Number of Reasoning Paths
-To further assess the contribution of reasoning path diversity, we conducted ablation experiments varying the number of generated paths $K\in\{2,3,4\}$ to evaluate their effects on execution accuracy, visualization quality and interpretability.
-Evaluation has been conducted on both the MatPlotBench and Qwen-Agent Code Interpreter benchmark datasets.
-The study reveals an interesting pattern observed when the number of reasoning path is reduced.
-As illustrated in Table 2 and Table 3, when $K=2$, the system exhibited limited diversity in which occasionally results in missing nuanced interpretations of complex user queries.
-Model K MatPlotBench
-Plot Score Executable Rate (%)
-GPT-4o mini 2 64.02 58
-3 66.12 60
-4 64.68 62
-Gemini 2.0 Flash 2 56.59 56
-3 59.37 61
-4 54.60 54
-Table 2: Results of ablation study on MatPlotBench.
-Model K Qwen-Agent
-Hard Easy Avg.
-GPT-4o mini 2 67.50 55.72 61.61
-3 70.68 57.23 63.96
-4 66.42 50.25 58.34
-Gemini 2.0 Flash 2 74.21 57.50 65.86
-3 80.79 57.17 68.98
-4 76.41 54.50 65.46
-Table 3: Results of ablation study on Qwen-Agent Code Interpreter benchmark.
-Conversely, increasing $K$ to 4 introduces extra paths which can sometimes add noise as less relevant interpretations are considered.
-Specifically, we observed cases where redundant or overly complex visualization components were generated, hence making the final output harder to be interpreted.
-While the additional reasoning paths increased the overall exploration space, they also required more extensive filtering and selection, leading to suboptimal execution efficiency.
-Notably, $K=3$ consistently achieves the best balance by providing sufficient diversity without introducing excessive noise.
-With too few reasoning paths, some nuanced aspects of ambiguous queries are lost, while too many paths configuration result in unnecessary exploration of low relevancy solutions.
-The number of reasoning path diversity set as $K=3$ optimally provides sufficient diversity without overcomplicating the reasoning process.
-4.3.2 Robustness with a Simple Integration Strategy
-To further validate the robustness of VisPath, we also evaluate an alternative integration strategy that simplicies the aggregation of multiple reasoning paths.
-In contrast to our full feedback-driven iterative approach which refines visualization code through multiple rounds of vision-language feedback, this streamlined method directly aggregates candidate outputs from different reasoning trajectories without intermediate corrections.
-Specifically, in this alternative strategy, three candidate codes, each derived from different reasoning paths, are directly combined to generate the final visualization output.
-This methodology reduces computational overhead and processing time while maintaining a degree of interpretability and correctness.
-Through examining the impact of this approach, we gain insight into the extent to which VisPath’s core strength stems from its multi-path reasoning capability versus its iterative refinement process.
+To investigate the contribution of reasoning path diversity, we conducted ablation experiments by varying the number of generated reasoning paths $K$.
+In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path multiplicity on the overall performance of VisPath, as shown in Figure 3.
+We observe a consistent pattern across all model and dataset combinations: performance improves as $K$ increases from 2 to 3, confirming that limited diversity $(K=2)$ often fails to capture nuanced interpretations of user queries.
+For example, with GPT-4o mini on MatPlotBench, the Plot Score improves from 64.02 to 66.12 (+2.10), and the Executable Rate improves from 58% to 60% (+2 points).
+On Gemini 2.0 Flash, the Plot Score increases from 56.59 to 59.37 (+2.78), and the Executable Rate from 56% to 61% (+5 points).
+While $K=4$ achieves the highest executable rate on MatPlotBench with GPT-4o mini (62%), we further extend our analysis up to $K$ = 8 to comprehensively assess the impact of reasoning path diversity.
+However, beyond $K=4$, we observe diminishing returns and even performance degradation, which is likely due to noisy or redundant reasoning paths.
+While added diversity initially aids interpretation, excessive expansion burdens the integration process and reduces overall efficiency.
+Among all configurations tested up to $K=8$, $K=3$ emerges as the most balanced choice, offering substantial performance gains in both the Executable Rate and the Plot Score while avoiding the inefficiencies observed at higher values of $K$.
+Hence, we adopt $K=3$ as the default configuration throughout our experiments.
+4.3.2 Robustness with a Simple Integration
+We evaluate an alternative integration strategy that simplifies the aggregation of multiple reasoning paths to further validate the robustness of VisPath.
 Model Feedback MatPlotBench Qwen-Agent
@@
 Table 2: Performance comparison of VisPath with and without visual feedback.
-The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two llm models.
-As shown in Table 4, even under this simplified integration, the performance of VisPath significantly outperforms all baseline methods, demonstrating its robustness and adaptability.
-This result suggests that the diversity of reasoning paths alone ensures a resilient and effective outcome, even in the absence of extensive intermediate corrections.
-This finding shows that the strength of our framework largely stems from its multi-path reasoning design.
-Even when the integration is simplified, the diversity of reasoning paths ensures that the aggregated output remains robust and effective, thus showing the overall adaptability and effectiveness of our approach.
+The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two LLMs.
+Instead of refining each candidate visualization with feedback from vision-language feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections.
+This setup reduces computational overhead and execution time while preserving the benefits of interpretive diversity.
+As shown in Table 2, even under this simplified configuration, VisPath outperforms all baseline methods, confirming that Multi-Path Reasoning alone offers a strong foundation for visualization code generation.
+While full feedback-driven optimization leads to additional performance improvements, this result highlights that the primary strength of VisPath lies in its capacity to explore and leverage diverse reasoning trajectories.
+The framework remains effective and adaptable, even with minimal refinements, further validating importance of its core design.
+4.3.3 Distinct Contribution of Visual Feedback
+To assess the role of visual feedback in improving code quality, we compare two variants of our framework.
+The first, VisPath (w/ feedback), uses a VLM to evaluate both rendered plots and error messages.
+The second, VisPath Execute (w/ binary feedback), simplifies evaluation by relying solely on the binary success or failure of code execution.
+Incorporating rendered visual feedback improves the Executable Rate by $2\%-4\%$ and consistently boosts the Plot Score across both LLMs, as shown in Table 3.
+On GPT-4o mini, Plot Score increases from 64.82 to 66.12 (+1.30) and Executable Rate from 58% to 60% (+2 points).
+On Gemini 2.0 Flash, Plot Score rises from 57.68 to 59.37 (+1.69), and Executable Rate from 59% to 63% (+4 points).
+Despite the numerical gains are modest, the results demonstrate the unique value of structured visual evaluation.
+Visual feedback enables more refined and user-aligned outputs by capturing subtle rendering issues that may not affect executability, demonstrating its importance in the final synthesis stage.
+LLM Variant Plot Score Executable Rate (%)
+GPT-4o mini (w) feedback 66.12 60
+(w) binary feedback 64.82 58
+Gemini 2.0 Flash (w) feedback 59.37 63
+(w) binary feedback 57.68 59
+Table 3: Ablation results isolating the impact of rendered visual feedback.
+Comparing the full VisPath model using structured plot-based feedback with a binary-only feedback variant.
+Visual feedback provides consistent gains in both Plot Score and Executable Rate across LLMs.
 5 Discussion
-Our main experiment and ablation studies decisively demonstrate that VisPath’s multi-path exploration revolutionizes visualization code generation by making it not just robust, but truly interpretable.
-[^3] Case study is provided in Appendix B.
-Unlike traditional methods that often yield simplistic, one-dimensional results, VisPath embraces complexity—delving deeper to capture nuances of user intent that would otherwise be lost.
-By integrating multiple perspectives during reasoning, our framework ensures that final visualizations are not only data-accurate but also rich in context, inherently intuitive, and unmistakably aligned with the user’s true objectives.
-The real breakthrough with VisPath is how it transforms visualization into a structured yet fluid process.
-Traditional methods often act like static translators, converting user queries into plots while neglecting the vital auxiliary components—legends, labels, line styles, and other elements—that enable true interpretability.
-They assume that a single pass can sufficiently capture the essence of a visualization, but this leads to visualizations that might technically be correct but fail at communication.
-In contrast, VisPath’s multi-path approach systematically explores different ways to represent data, integrating feedback at each step.
-This ensures that even subtle yet critical design elements are not just included, but optimized, making the resulting visualizations feel complete, polished, and immediately comprehensible.
-Beyond accuracy, VisPath introduces a level of adaptability that is indispensable in today’s fast-evolving data landscape.
-It’s not just a tool for static reports—it is built for real-time streaming data, dynamic dashboards, and interactive analytics.
-Even with broad or loosely defined prompts, our framework intelligently synthesizes high-quality visualizations, closing the gap between intent and execution.
-As industries continue to demand clearer, faster, and more adaptable insights, VisPath is poised to lead the way, setting a new standard for intelligent, context-aware data visualization.
+Our proposed VisPath framework substantially advances visualization code generation by addressing the core weaknesses of existing methods: limited interpretive flexibility and insufficient refinement.
+By employing Multi-Path Reasoning, VisPath explores diverse interpretations of user intent, which leads to more accurate visualizations, especially for ambiguous queries.
+Experimental results confirm its superiority: VisPath outperforms all baselines on both MatPlotBench and the Qwen-Agent Code Interpreter benchmark, with up to 9.14% improvement in Plot Score and 10% in Executable Rate.
+Ablation studies further validate VisPath’s design.
+First, increasing the number of reasoning paths enhances both visual quality and code executability.
+Second, even without visual feedback, Multi-Path Reasoning alone proves highly effective.
+Third, using structured plot-based feedback, rather than binary execution signals, significantly improves output alignment with user intent, confirming the value of our feedback-driven optimization module.
+The figures provided in Appendix clearly illustrate VisPath’s robustness.
+[^4] The detailed cases are provided in Appendix B.
+In Case 1, it correctly centers conditional means using adjusted strip widths, unlike other methods that misalign pointplots or ignore key parameters.
+In Case 2, VisPath generates a proper polar bar chart with radial coordinates and legible labels, while others either fail to render or produce standard rectangular plots, with text often overlapping or unreadable.
+Case 3 highlights VisPath’s ability to handle compositional visualization requests: it correctly interprets the instruction to "visualize in 3 different ways" by generating a multi-subplot layout that includes line, scatter, and bar plots.
+Other approaches, by contrast, either collapsed all time series into a single chaotic plot or failed to differentiate the visualization modalities at all.
+These examples demonstrate how VisPath’s Multi-Path Reasoning and visual feedback lead to more precise and semantically aligned visualizations than current alternatives.
 6 Conclusion
-In this work, we present VisPath, a groundbreaking framework that redefines the landscape of automated visualization code generation over existing methods.
-Unlike prior methods, our approach seaminglessly combines multi-path reasoning with feedback-driven optimization.
-Accurately capturing diverse user intents and refine generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark.
-By prioritizing adapability and user intent alignment, VisPath is uniquely positioned to handle compledities of real-world data visualization tasks.
-Looking ahead, future work could explore VisPath’s adaptability in more dynamic, real-world visualization scenarios, further broadening its scope and enhancing its practical utility in complex data analysis contexts.
+In this work, we present VisPath, a framework that leverages Multi-Path Reasoning and feedback-driven optimization to enhance automated visualization code generation.
+Unlike prior methods, our approach seamlessly combines Multi-Path Reasoning with feedback-driven optimization.
+Accurately capturing diverse user intents and refining generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark.
+By prioritizing adaptability, VisPath is uniquely positioned to handle ambiguous user queries through a combination of diverse reasoning paths and visual feedback integration.
+Future work could explore VisPath’s adaptability in more dynamic, real-world scenarios, further broadening its scope and practical utility in complex data analysis contexts.
 7 Limitations
-Despite its effectiveness, the current iteration of our framework relies on a limited feedback mechanism, evaluating visualizations primarily based on two aspects: query-code alignment and query-plot image alignment.
-While these aspects provide valuable insights, they may not fully capture the granular elements that contribute to overall interpretability.
-To enhance the depth of feedback itself, future research could focus on decomposition of assessment process, assessing individual plot components separately.
-This granular analysis would render for a more nuanced assessment of readability, element appropriateness and visual coherence, ultimately leading to more refined visualization code generation.
-Strengthening the feedback mechanism in such way will be crucial for maximizing VisPath’s effectiveness in diverse and complex visualization scenarios.
+Despite its effectiveness, the current framework relies on a limited feedback mechanism focused on query-code and query-plot alignment.
+While informative, these signals may overlook finer-grained elements essential to interpretability.
+Thus, future work could improve feedback depth by assessing individual plot components, such as readability and visual coherence, enabling more precise and refined visualization code generation.
+Moreover, while achieving strong performance, VisPath requires several rounds of agent interaction, including multi-path reasoning, execution, and feedback integration, which may introduce inefficiencies in certain use cases.
+Future work could explore ways to selectively identify the most promising reasoning paths early in the process, reducing redundant computation while preserving the benefits of diverse interpretation.
 Appendix A Appendix A. Prompts Used
@@
 Ensure that you first analyze the given data description and create queries that align with the data.
-If no data description is provided, follow the original query as is.
+If no data description is provided, follow the initial query as is.
 You must follow the Python list format for the output.
 Do not modify the detailed instructions from the original user query.
-Original query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
+initial query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
 Prompt for Code Generation [System Prompt] You are an expert in data visualization code generation.
@@
 Appendix B Appendix B. Case Study
-Figure 3: Overview of end-to-end process of visualizing women’s million-dollar earnings data using VisPath.
-It includes the user query, raw data, multi-path reasoning, generated figures, visual feedback, final script, and the completed visualization.
+Figure 4: Overview of the end-to-end process of visualizing conditional means alignment.
+Figure 5: Overview of the end-to-end process of generating a polar bar chart.
+Figure 6: Overview of the end-to-end process of visualizing time series data in three different ways.

### v3

@@
-Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
+
 Abstract
-Rapid advancements in Large Language Models (LLMs) have accelerated their integration into automated visualization code generation applications.
-Despite advancements through few-shot prompting and query expansion, existing methods remain limited in handling ambiguous and complex queries, thereby requiring manual intervention.
+Large Language Models (LLMs) have become a cornerstone for automated visualization code generation, enabling users to create charts through natural language instructions.
+Despite improvements from techniques like few-shot prompting and query expansion, existing methods often struggle when requests are underspecified in actionable details (e.g., data preprocessing assumptions, solver or library choices, etc.), frequently necessitating manual intervention.
 To overcome these limitations, we propose VisPath: a Multi-Path Reasoning and Feedback-Driven Optimization Framework for Visualization Code Generation.
 VisPath handles underspecified queries through structured, multi-stage processing.
-It begins by reformulating the user input via Chain-of-Thought (CoT) prompting, which refers to the initial query while generating multiple extended queries in parallel, enabling the LLM to capture diverse interpretations of the user intent.
+It begins by using Chain-of-Thought (CoT) prompting to reformulate the initial user input, generating multiple extended queries in parallel to surface alternative plausible concretizations of the request.
 These queries then generate candidate visualization scripts, which are executed to produce diverse images.
 By assessing the visual quality and correctness of each output, VisPath generates targeted feedback that is aggregated to synthesize an optimal final result.
-Extensive experiments on widely-used benchmarks including MatPlotBench and the Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, offering a more reliable solution for AI-driven visualization code generation.
-Automated Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization
-Wonduk Seo1,2   Seungyong Lee   Daye Kang1,3 Hyunjin An1   Zonghao Yuan4   Seunghyun Lee1thanks: Corresponding author.
-1Enhans   2Peking University   3KAIST   4Tsinghua University 1{wonduk,daye,hyunjin,seunghyun}@enhans.ai
+Extensive experiments on MatPlotBench and Qwen-Agent Code Interpreter Benchmark show that VisPath outperforms state-of-the-art methods, providing a more reliable framework for AI-driven visualization generation.
 1 Introduction
-Figure 1: Overview of different approaches for visualization code generation.
-Comparing two baseline methods, namely Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al.
-(2024), with our proposed VisPath framework.
-Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets Vondrick et al.
-(2013); Demiralp et al.
-(2017); Unwin (2020); Li et al.
-(2024a).
-Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js Barrett et al.
-(2005); Bisong and Bisong (2019); Zhu (2013).
-These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users Bresciani and Eppler (2015); Saket et al.
-(2018); Sharif et al.
-(2024).
-As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible Wang et al.
-(2015); Dibia and Demiralp (2019); Qian et al.
-(2021).
-In response to this challenge, Large Language Models (LLMs) have emerged as a promising solution for simplifying visualization creation Wang et al.
-(2023a); Han et al.
-(2023); Xie et al.
-(2024).
-By translating natural language instructions into executable code, LLM-based systems eliminate the need for extensive programming knowledge, allowing users to generate visualizations more intuitively Xiao et al.
-(2023); Ge et al.
-(2023); Zhang et al.
-(2024b).
-Recent visualization methods such as ChartGPT Tian et al.
-(2024) and NL4DV Sah et al.
-(2024) demonstrate the potential of LLMs to provide interactive, conversational interfaces for visualization.
-These systems enable users to create complex charts with minimal effort, bridging the gap between technical expertise and effective data exploration Dibia (2023); Kim et al.
-(2024).
-More recently, LLM-based visualization frameworks such as Chat2VIS Maddigan and Susnjak (2023) and MatPlotAgent Yang et al.
-(2024) have been introduced to improve automated visualization code generation.
+Data visualization has long been an essential tool in data analysis and scientific research, enabling users to uncover patterns and relationships in complex datasets [28, 9, 27, 14].
+Traditionally, creating visualizations requires manually writing code using libraries such as Matplotlib, Seaborn, or D3.js [2, 3, 40].
+These approaches demand programming expertise and significant effort to craft effective visual representations, which can be a barrier for many users [4, 23, 25].
+As datasets continue to grow in size and complexity, researchers have explored ways to automate visualization generation, aiming to make the process more efficient and accessible [31, 10, 21].
+Large Language Models (LLMs) have emerged as a promising solution for simplifying visualization creation [29, 13, 36, 32].
+By translating natural language instructions into executable code, LLM-based systems eliminate the need for extensive programming knowledge, allowing users to generate visualizations more intuitively [35, 11, 39].
+More recently, Chat2VIS [19] and MatPlotAgent [37] have been introduced to improve automated visualization code generation.
 Specifically, Chat2VIS follows a prefix-based approach, guiding LLMs to generate visualization code consistently; and MatPlotAgent expands the query before code generation.
-However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions and unable to fix-out when caught in misleading bugs; (2) they rely on predefined structures or examples which restrict adaptability to ambiguous or unconventional user queries; (3) they encapsulate limitation in their inability to aggregate and synthesize multi-dimensional feedback.
+However, these methods face several limitations: (1) they generate code in a single-path manner, limiting exploration of alternative solutions, and are unable to recover when generating erroneous code; (2) they rely on predefined structures or examples, which restrict adaptability to ambiguous or unconventional user queries; and (3) a fundamental limitation is their inability to aggregate and synthesize multi-dimensional feedback.
 Without a mechanism to retrieve outputs that reflect diverse possibilities, they face difficulties in capturing the intricate details required for visualizations that are both functionally precise and contextually relevant.
-To address these limitations, we introduce VisPath: A Branch Exploration Framework for Visualization Code Synthesis via Multi-Path Reasoning and Feedback-Driven Optimization, a novel approach that redefines how visualization code is generated, as illustrated in Figure 1.
-Traditional methods often fall short of delivering the depth and precision users truly need, struggle to capture the intricate details that make a visualization not just functional but meaningful.
-VisPath challenges this limitation by incorporating Multi-Path Reasoning and Feedback-Driven Optimization, systematically exploring multiple interpretative pathways to construct a more accurate, context-aware, and fully executable visualization.
-Rather than simply translating user input into code, VisPath ensures that every critical aspect, both explicitly stated and implicitly necessary, is carefully considered, creating a visualization that is not only correct but insightful.
-At its core, it generates multiple reasoning paths that analyze the user’s intent from different perspectives, producing structured blueprints that are then transformed into visualization scripts through Chain-of-Thought (CoT) prompting.
-These multiple candidates are evaluated using a Vision-Language Model (VLM) to assess accuracy, clarity, and alignment with the intended message.
-The results are then refined through an Aggregation Module, optimizing the final output for both reliability and impact.
-Extensive experiments on benchmark datasets, including MatPlotBench Yang et al.
-(2024) and Qwen-Agent Code Interpreter Benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md, demonstrate that VisPath notably outperforms state-of-the-art methods across all evaluation metrics.
-In addition, ablation studies confirm that VisPath’s performance gains primarily stem from its Multi-Path Reasoning mechanism and feedback-driven optimization.
-By systematically generating and evaluating multiple reasoning paths and leveraging feedback aggregation, VisPath significantly enhances the accuracy, robustness against underspecified queries, and adaptability to diverse user intents.
-Our investigations demonstrate its ability to capture nuanced user intents, improve execution reliability, and minimize errors, making visualization code generation more accessible and effective for domains such as business intelligence, scientific research, and automated reporting.
+To address underspecified visualization requests, we propose VisPath, a Multi-Path Reasoning and Feedback-Driven Optimization framework.
+VisPath samples multiple plausible concretizations of implicit implementation choices, executes candidate scripts, and leverages visual feedback to synthesize a robust final program.
+Rather than directly translating user input into code, VisPath systematically accounts for both explicit requirements and implicit necessities to produce visualizations that are correct and insightful.
+It generates multiple reasoning paths that interpret user intent from different perspectives, producing structured blueprints that are converted into visualization scripts via Chain-of-Thought prompting.
+These candidates are then evaluated by a Vision-Language Model (VLM) for accuracy, clarity, and alignment with the intended message, and refined by a synthesis agent to optimize reliability and impact.
+Experiments show that VisPath improves plot-level correctness and executability over prompting and agent-based baselines.
+Ablations attribute the gains to multi-path exploration and feedback integration, demonstrating stronger intent capture, higher execution reliability, and fewer errors—making visualization code generation more accessible for business intelligence, scientific research, and automated reporting.
 2 Related Work
-Numerous methods have been applied for Text-to-Visualization (Text2Vis) generation, which has significantly evolved over the years, adapting to new paradigms in data visualization and natural language processing Dibia and Demiralp (2019); Wu et al.
-(2022); Chen et al.
-(2022a, b); Rashid et al.
-(2022); Zhang et al.
-(2024a).
-Early approaches such as Voyager Wongsuphasawat et al.
-(2015) and Eviza Setlur et al.
-(2016) largely relied on rule-based systems, which mapped textual commands to predefined chart templates or specifications through handcrafted heuristics de Araújo Lima and Diniz Junqueira Barbosa (2020).
-While these methods demonstrated the feasibility of automatically converting text into visualizations Moritz et al.
-(2018); Cui et al.
-(2019), they often required extensive domain knowledge and struggled with more nuanced or ambiguous user requirements Li et al.
-(2021); Wang et al.
-(2023b).
-Inspired by developments in deep learning, researchers began to incorporate neural networks to handle free-form natural language and broaden the range of supported visualization types Liu et al.
-(2021); Luo et al.
-(2021).
-Figure 2: Overview of the proposed VisPath framework for creating robust visualization code generation.
-The framework consists of combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
+Numerous methods have been applied for Text-to-Visualization (Text2Vis) generation, which has significantly evolved over the years, adapting to new paradigms in data visualization and natural language processing [10, 34, 5, 6, 22, 38].
+Early approaches such as Voyager [33] and Eviza [24] largely relied on rule-based systems, which mapped textual commands to predefined chart templates or specifications through handcrafted heuristics [8].
+While these methods demonstrated the feasibility of automatically converting text into visualizations [20, 7], they often required extensive domain knowledge and struggled with more nuanced or ambiguous user requirements [15, 30].
+Inspired by developments in deep learning, researchers began to incorporate neural networks to handle free-form natural language and broaden the range of supported visualization types [17, 18].
 Building on these machine learning strategies, numerous studies have utilized LLMs to further enhance system flexibility.
-Recent frameworks such as Chat2VIS Maddigan and Susnjak (2023) and Prompt4Vis Li et al.
-(2024b) utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting.
-More recent approaches, such as MatPlotAgent Yang et al.
-(2024) and PlotGen Goswami et al.
-(2025), extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations.
+Recent frameworks such as Chat2VIS [19] and Prompt4Vis [16] utilize few-shot learning or query expansion to refine user queries, subsequently generating Python visualization scripts through instruction-based prompting.
+More recent approaches, such as MatPlotAgent [37] and PlotGen [12], extend these frameworks by integrating a vision-language feedback model to iteratively optimize the final code based on evaluations of the rendered visualizations.
 The aforementioned approaches often struggle to effectively capture user intent in complex visualization tasks.
@@
 3 Methodology
-We introduce VisPath, a framework for robust visualization code generation that leverages diverse reasoning and visual feedback.
-VisPath is built on three core components: (1) Multi-Path Query Expansion, which generates multiple reasoning paths informed by the dataset description; (2) Code Generation from Expanded Queries, which synthesizes candidate visualization scripts via Chain-of-Thought (CoT) prompting while grounding them in the actual data context; and (3) Feedback-Driven Code Optimization, where a VLM evaluates and refines the outputs to ensure generation robust visualization code.
-An overview of this process is shown in Figure 2.
+We introduce VisPath, a framework for robust visualization code generation that combines diverse reasoning with visual feedback.
+Given a user query $Q$ and a dataset description $D$, VisPath proceeds in three stages: (i) multi-path query expansion, (ii) visualization code synthesis and execution, and (iii) feedback-driven refinement.
+An overview is shown in Figure 1.
+Figure 1: Overview of the proposed VisPath framework for creating robust visualization code generation.
+The framework consists of a combination of Multi-Path Agent, Visual Feedback Agent, and Synthesis Agent.
+Notation.
+We denote any agent component by $G_{\bullet}(\,\cdot\mid S_{\bullet})$, where $S_{\bullet}$ is the system prompt that specifies the role, constraints, and output format for that component.
+In contrast, code execution is performed by a non-Agent operator, denoted $\mathcal{E}(\cdot)$.
 3.1 Multi-Path Generation
-The potential for rigid interpretation is a major limitation in visualization code generation, a single query can have multiple valid representations depending on its dataset structure.
-VisPath mitigates this by generating multiple extended queries within a single interaction.
-Given a user query $Q$ and a corresponding dataset description $D$, a Multi-Path Agent is employed to expand the query into $K$ distinct reasoning pathways:
-$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}=f_{\_}{\text{mpa}}(Q,D),$ (1)
-where $f_{\_}{\text{mpa}}$ denotes the function of the Multi-Path Agent implemented via an LLM.
-The dataset description $D$ plays a crucial core in shaping these reasoning paths by providing contextual information about variable types, inherent relationships, and the suitability of different chart types for a more grounded interpretation of the query.
-Each $R_{\_}i$ serves as a detailed logical blueprint outlining one possible approach to fulfill the visualization request.
-The purpose of generating multiple reasoning paths is not simply to increase their quantity, but to ensure interpretive diversity.
-Vispath mitigates the risk of depending on a single, potentially incorrect assumption by intentionally exploring alternative and context-aware reasoning strategies.
-This design ensures that our framework effectively considers a broad range of potential interpretations, thereby increasing the quality of reasoning and the likelihood of capturing the true user intent even when queries are ambiguous or underspecified.
-3.2 Code Generation from Reasoning Paths
-Once diverse reasoning paths are established, the subsequent stage involves translating each path into executable Python scripts.
-For each reasoning path $R_{\_}i$ generated in Equation 1, a dedicated Code Generation LLM produces the corresponding visualization code using Chain-of-Thought (CoT) prompting:
-$C_{\_}i=f_{\_}{\text{code}}(D,R_{\_}i),$ (2)
-where $f_{\_}{\text{code}}$ represents the code generation function.
-The dataset description $D$ is explicitly provided to ground the generated code in the actual data context, ensuring that variable names, data types, and visualization parameters align correctly with the underlying data attributes.
-The generated code $C_{\_}i$ is then executed to render a candidate visualization:
-$V_{\_}i=f_{\_}{\text{exec}}(C_{\_}i),\quad i=1,2,\dots,K,$ (3)
-with $f_{\_}{\text{exec}}$ serving as the code execution function.
-By executing the code directly, we ensure that the visualization accurately reflects the intended operations without reintroducing the dataset description $D$ at this stage, as the code has already been conditioned on $D$ during its generation.
-In practice, some generated codes may not be executable. Rather than engaging in an explicit debugging loop, we record the execution status as a binary executability indicator:
-$\epsilon_{\_}i=\begin{cases}1,&\text{if }C_{\_}i\text{ is executable}\\ 0,&\text{otherwise}.\end{cases}$ (4)
-To route the outputs appropriately, we introduce:
-$Z_{\_}i=\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1,\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0.\end{cases}$ (5)
-The result $Z_{\_}i$ (either the rendered visualization or the error message) is then provided, along with $C_{\_}i$ and the initial query $Q$, to the feedback model in the subsequent stage.
+A single visualization request can support multiple valid interpretations depending on dataset schema, variable types, and implicit relationships.
+To mitigate brittle, single-assumption interpretations, VisPath generates a diverse set of reasoning paths conditioned on the dataset context.
+Specifically, a Multi-Path Agent produces $K$ distinct reasoning pathways:
+$\{R_{1},R_{2},\dots,R_{K}\}\sim G_{\text{mpa}}(Q,D\mid S_{\text{mpa}}).$ (1)
+Here each $R_{i}$ is a structured plan (a logical blueprint) describing one plausible data-grounded interpretation of $Q$ under $D$ (e.g., chart type choice, grouping/aggregation assumptions, encoding decisions, and potential preprocessing).
+Conditioning on $D$ encourages paths that respect variable names, data types, and feasible visual encodings.
+3.2 Code Generation and Execution
+For each reasoning path $R_{i}$, a code-generation agent synthesizes an executable Python visualization script:
+$C_{i}\sim G_{\text{code}}(D,R_{i}\mid S_{\text{code}}),\quad i=1,\dots,K.$ (2)
+The dataset description $D$ is provided explicitly to ground the script in the actual data context (e.g., column names, semantic types, and plausible transformations), reducing hallucinated variables and mismatched encodings.
+Each generated script is then executed by a deterministic execution operator $\mathcal{E}$.
+We define the routed execution output as a tuple that jointly encodes executability and the observable outcome:
+$Z_{i}:=\mathcal{E}(C_{i})=\begin{cases}\bigl(1,\,\mathrm{plot}(V_{i})\bigr),&\text{if }C_{i}\text{ executes successfully and renders }V_{i},\\ \bigl(0,\,\mathrm{err}(C_{i})\bigr),&\text{otherwise},\end{cases}$ (3)
+for $i=1,\dots,K$, where the first element indicates executability and the second element is either the rendered plot image or the execution error message.
+This design avoids an explicit debugging loop while still preserving informative failure signals for downstream refinement.
 3.3 Feedback-Driven Code Optimization
-While most code generation frameworks primarily focus on producing syntactically correct scripts, our framework extends beyond this by incorporating an additional mechanism.
-As final stage, VisPath synthesizes the most robust and accurate visualization code by leveraging both the executability information and structured visual feedback.
-A VLM is employed to analyze each candidate by evaluating the initial query $Q$, the generated code $C_{\_}i$, and the routed output $Z_{\_}i$.
-This evaluation is formalized as:
-$F_{\_}i=f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i),$ (6)
-where $F_{\_}i$ provides structured feedback on key aspects such as chart layout, the alignment between the intended request and the rendered visualization (or error context), and visual readability (including potential improvements).
-To capture the complete quality signal from each candidate, we pair the feedback with its corresponding generated code:
-$S_{\_}i=(C_{\_}i,F_{\_}i),\quad i=1,2,\dots,K.$ (7)
-Leveraging the collective code-feedback pairs along with the initial query $Q$ and the dataset description $D$, an Integration Module synthesizes the final, refined visualization code:
-$C^{*}=f_{\_}{\text{integrate}}\left(Q,D,\{S_{\_}i\}_{\_}{i=1}^{K}\right),$ (8)
-where $C^{*}$ represents the optimized visualization code and $f_{\_}{\text{integrate}}$ denotes the function that aggregates the strengths of each candidate code alongside its corresponding feedback.
-This formulation ensures that the final code is not only constructed based on the insights extracted from the candidate outputs but is also meticulously aligned with the original user query and the provided dataset description.
-An algorithm for VisPath is detailed in Algorithm 1 as shown below:
-Algorithm 1 Algorithm for VisPath
-1:User query $Q$, dataset $D$, number of reasoning paths $K$, Multi-Path Agent $f_{\_}{\text{mpa}}$, Code Generation LLM $f_{\_}{\text{code}}$, Code Execution Function $f_{\_}{\text{exec}}$, Feedback Model $f_{\_}{\text{feedback}}$, Integration Module $f_{\_}{\text{integrate}}$
-2:// Step 1: Multi-Path Query Expansion
-3:$\{R_{\_}1,R_{\_}2,\dots,R_{\_}K\}\leftarrow f_{\_}{\text{mpa}}(Q,D)$
-4:// Step 2: Code Generation from Reasoning Paths
-5:for $i=1$ to $K$ do
-6:     $C_{\_}i\leftarrow f_{\_}{\text{code}}(D,R_{\_}i)$
-7:     $V_{\_}i\leftarrow f_{\_}{\text{exec}}(C_{\_}i)$
-8:     $\epsilon_{\_}i\leftarrow\begin{cases}1,&\text{if }C_{\_}i\text{ executes successfully}\\ 0,&\text{otherwise}\end{cases}$
-9:     $Z_{\_}i\leftarrow\begin{cases}\text{plot image}(V_{\_}i),&\text{if }\epsilon_{\_}i=1\\ \text{error message from }C_{\_}i,&\text{if }\epsilon_{\_}i=0\end{cases}$
-10:end for
-11:// Step 3: Feedback-Driven Code Optimization
-12:for $i=1$ to $K$ do
-13:     $F_{\_}i\leftarrow f_{\_}{\text{feedback}}(Q,C_{\_}i,Z_{\_}i)$
-14:     $S_{\_}i\leftarrow(C_{\_}i,F_{\_}i)$
-15:end for
-16:$C^{*}\leftarrow f_{\_}{\text{integrate}}\bigl{(}Q,D,\{S_{\_}1,S_{\_}2,\dots,S_{\_}K\}\bigr{)}$
-17:return $C^{*}$
+Next, a feedback agent evaluates each candidate by jointly considering the user intent and the observed execution outcome:
+$F_{i}\sim G_{\text{fb}}(Q,C_{i},Z_{i}\mid S_{\text{fb}}),\quad i=1,\dots,K,$ (4)
+where $F_{i}$ is structured feedback capturing (i) semantic alignment with $Q$, (ii) correctness with respect to the data context implied by $D$, and (iii) visual quality/readability (e.g., labeling, legends, scales, occlusion, layout).
+Finally, a Synthesis Agent synthesizes a refined visualization program by aggregating the full set of candidates and their feedback:
+$C^{*}\sim G_{\text{syn}}\!\left(Q,D,\{(C_{i},F_{i})\}_{i=1}^{K}\mid S_{\text{syn}}\right).$ (5)
+The output $C^{*}$ is optimized to be (1) executable, (2) faithful to the user intent in $Q$, and (3) visually informative given the dataset context $D$, by selectively inheriting strengths and correcting weaknesses identified across the candidate set.
+Summary.
+VisPath generates multiple dataset-aware reasoning paths, translates them into candidate scripts, executes each script to obtain either a plot or an error signal, and then uses structured multimodal feedback to synthesize these signals into a final robust visualization program.
 4 Experiments
@@
 4.1.1 Experimental Datasets
-We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench Yang et al.
-(2024) and Qwen-Agent Code Interpreter Benchmark.
+We evaluate our approach on two Text-to-Visualization benchmarks: MatPlotBench [37] and Qwen-Agent Code Interpreter Benchmark.
 Specifically, MatPlotBench comprises 100 items with ground truth images; we focus on its simple instruction subset for nuanced queries.
-In contrast, the Qwen-Agent Code Interpreter Benchmark includes 295 records: 163 related to visualization, and evaluates Python code interpreters on tasks such as mathematical problem solving, data visualization, and file handling based on Code Executability and Code Correctness.
+In this paper, we use underspecification to mean that a query omits at least one decision that must be instantiated in code (e.g., plot construction primitives, data transformations, or layout conventions).
+Although each benchmark item provides a single reference visualization, many prompts still require such implicit choices.
+For example, a broken-axis plot typically entails a two-panel shared-x layout with different y-limits, and polar bar charts require a polar projection with angles in radians.
+VisPath is designed to surface and resolve these implicit choices via multi-path concretization.
+The Qwen-Agent Code Interpreter Benchmark contains 295 records, of which 163 are visualization-related, and evaluates Python code-interpreter agents using Code Executability and Code Correctness on tasks including data visualization.
 4.1.2 Models Used
-Large Language Models (LLMs):
-For the code inference stage, we experiment with GPT-4o mini Achiam et al.
-(2023) and Gemini 2.0 Flash Team et al.
-(2024) to generate candidate visualization code from the reasoning paths.
-Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work Yang et al.
-(2024).
-To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o Achiam et al.
-(2023) and Gemini 2.0 Flash Team et al.
-(2024) as our visualization feedback model, which provides high-quality reference assessments.
-Vision-Language Models (VLMs):
+Large Language Models (LLMs)
+For the code inference stage, we experiment with GPT-4o mini [1] and Gemini 2.0 Flash [26] to generate candidate visualization code from the reasoning paths.
+Both models are configured with a temperature of 0.2 to ensure precise and focused outputs, in line with previous work [37].
+To evaluate the generated code quality and guide the subsequent optimization process, we utilize GPT-4o [1] and Gemini 2.0 Flash [26] as our visualization feedback model, which provides high-quality reference assessments.
+Vision-Language Models (VLMs)
 In order to assess the visual quality and correctness of the rendered plots, we incorporate vision evaluation models into our framework.
-Specifically, GPT-4o Achiam et al.
-(2023) is employed for detailed plot evaluation in all evaluation tasks.
+Specifically, GPT-4o [1] is employed for detailed plot evaluation in all evaluation tasks.
 This setup ensures the thorough evaluation of both the syntactic correctness of the code and the aesthetic quality of the resulting visualizations.
@@
 In our experiments, we utilized evaluation metrics introduced by previous work to ensure consistency and comparability.
-MatPlotBench Bisong and Bisong (2019) assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions.
-Qwen-Agent Code Interpreter benchmark [^2] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/codeinterpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy.
+MatPlotBench [3] assesses graph generation models using two key metrics: Plot Score, which measures similarity to the Ground Truth (0–100), and Executable Score, which represents the percentage of error-free code executions.
+Qwen-Agent Code Interpreter benchmark [^1] https://github.com/QwenLM/Qwen-Agent/blob/main/benchmark/code_interpreter/README.md evaluates visualization models based on Visualization-Hard and Visualization-Easy.
 Compared to MatPlotBench, Qwen-Agent Code Interpreter benchmark assesses image alignment via a code correctness metric.
-Previous studies showed that GPT-based VLM evaluations align well with human assessments Yang et al.
-(2024), hence VLM was used for evaluation.
+Previous studies showed that GPT-based VLM evaluations align well with human assessments [37], hence VLM was used for evaluation.
+4.1.4 Baseline Methods
+We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS [19] employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent [37] first expands the query and then refines the code via a self-debugging loop with feedback.
+Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output.
+For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses a critique-based debugging loop.
+Table 1: Performance comparison of various methods across different benchmarks.
+Visualization-Hard and Visualization-Easy refer to the Accuracy of Code Execution Results on different subsets of the Qwen-Agent Code Interpreter benchmark.
+Bold text indicates the best performance, underlined text indicates the second-best performance.
+denotes our proposed method.
 Model Methods MatPlotBench Qwen-Agent Code Interpreter benchmark
@@
 GPT-4o mini Zero-Shot 62.38 53 59.68 45.50 52.59
-CoT Prompting Wei et al. (2022) 61.95 50 57.50 40.00 48.75
-Chat2VIS Maddigan and Susnjak (2023) 56.98 53 59.36 36.50 47.93
-MatPlotAgent Yang et al. (2024) 63.90 58 67.50 53.25 60.38
+CoT Prompting 61.95 50 57.50 40.00 48.75
+Chat2VIS 56.98 53 59.36 36.50 47.93
+MatPlotAgent 63.90 58 67.50 53.25 60.38
 VisPath (Ours) 66.12 60 70.68 57.23 63.96
 Gemini 2.0 Flash Zero-Shot 55.00 54 68.97 52.18 60.58
-CoT Prompting Wei et al. (2022) 53.56 61 40.00 63.89 51.95
-Chat2VIS Maddigan and Susnjak (2023) 54.89 55 59.36 56.50 57.93
-MatPlotAgent Yang et al. (2024) 56.31 58 77.62 51.50 64.56
+CoT Prompting 53.56 61 40.00 63.89 51.95
+Chat2VIS 54.89 55 59.36 56.50 57.93
+MatPlotAgent 56.31 58 77.62 51.50 64.56
 VisPath (Ours) 59.37 63 80.79 57.17 68.98
-Table 1: Performance comparison of various methods across different benchmarks.
-Zero-Shot refers to directly generating code.
-CoT Prompting utilizes Chain of Thought Prompting.
-Visualization-Hard and Visualization-Easy refer to the Accuracy of Code Execution Results on different subsets of the Qwen-Agent Code Interpreter benchmark.
-Bold text indicates the best performance, underlined text indicates the second-best performance.
-denotes our proposed method.
-4.1.4 Baseline Methods
-We compare VisPath against competitive baselines: (1) Zero-Shot directly generates visualization code without intermediate reasoning, (2) CoT Prompting uses Chain-of-Thought (CoT) prompting to articulate its reasoning, while (3) Chat2VIS Maddigan and Susnjak (2023) employs guiding prefixes to mitigate ambiguity, and (4) MatPlotAgent Yang et al.
-(2024) first expands the query and then refines the code via a self-debugging loop with feedback.
-Moreover, our proposed framework VisPath generates three reasoning paths with corresponding visual feedback to refine the final output.
-[^3] Prompts are detailed in Appendix A.
-For a fair comparison aligned with our experimental setting, MatPlotAgent is limited to three iterations, and uses critique-based debugging loop as well.
 4.2 Experimental Analysis
-VisPath is evaluated against four baselines: Zero Shot prompting, CoT prompting, Chat2VIS, and MatPlotAgent.
-The evaluation is conducted on MatPlotBench and the Qwen-Agent Code Interpreter benchmark using GPT-4o mini and Gemini 2.0 Flash, as shown in Table 1.
+VisPath is evaluated against four baselines, as shown in Table 1.
 Zero-shot prompting generates visualization code directly from natural language queries without intermediate reasoning.
-While computationally efficient, it often struggles to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs.
+Despite its computational efficiency, this method frequently fails to handle ambiguity or under-specification, resulting in misaligned or incomplete outputs.
 On MatPlotBench (GPT-4o mini), it achieves a Plot Score of 62.38 and an Executable Rate of 53%.
@@
 Chat2VIS extends CoT prompting by adopting prefix templates to improve coherence and reduce ambiguity in user instructions.
-While this approach is effective for well-structured or common query formats, its dependence on fixed templates limits adaptability when processing loosely specified or novel queries.
+While effective for well-structured queries, its reliance on fixed templates limits adaptability to underspecified or unconventional requests.
 Such limitation is evident in its performance on MatPlotBench, where it achieves a Plot Score of 56.98 and an Executable Rate of 53%.
@@
 Evaluated across multiple benchmark settings, VisPath notably outperforms baselines, achieving up to 9.14 point gains in Plot Score and a 10% point increase in Executable Rate.
-These improvements well demonstrate VisPath’s robustness in exploring diverse reasoning paths and iteratively refine outputs through structured visual feedback, effectively reducing semantic ambiguity and improving execution reliability.
+These improvements demonstrate VisPath’s robustness in exploring diverse reasoning paths and refining outputs through visual feedback, which reduces semantic ambiguity and enhances execution reliability.
 4.3 Ablation Study
-To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments in this section.
+To further examine the robustness and design choices of VisPath, we conduct a series of ablation experiments.
 Specifically, we analyze the following three aspects: (i) varying the number of generated reasoning paths, (ii) the effect of removing visual feedback during integration, and (iii) the contribution of visual feedback beyond binary executability.
@@
 To investigate the contribution of reasoning path diversity, we conducted ablation experiments by varying the number of generated reasoning paths $K$.
-In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path multiplicity on the overall performance of VisPath, as shown in Figure 3.
+In particular, we extended the range of $K$ from 2 to 8 to examine the effect of increased path on the overall performance of VisPath, as shown in Figure 2.
 We observe a consistent pattern across all model and dataset combinations: performance improves as $K$ increases from 2 to 3, confirming that limited diversity $(K=2)$ often fails to capture nuanced interpretations of user queries.
-For example, with GPT-4o mini on MatPlotBench, the Plot Score improves from 64.02 to 66.12 (+2.10), and the Executable Rate improves from 58% to 60% (+2 points).
-On Gemini 2.0 Flash, the Plot Score increases from 56.59 to 59.37 (+2.78), and the Executable Rate from 56% to 61% (+5 points).
 While $K=4$ achieves the highest executable rate on MatPlotBench with GPT-4o mini (62%), we further extend our analysis up to $K$ = 8 to comprehensively assess the impact of reasoning path diversity.
@@
 We evaluate an alternative integration strategy that simplifies the aggregation of multiple reasoning paths to further validate the robustness of VisPath.
-Model Feedback MatPlotBench Qwen-Agent
-Plot Score Executable Rate (%) Avg.
-GPT-4o mini (w/o) feedback 63.76 56 58.00
-(w) feedback 66.12 60 63.96
-Gemini 2.0 Flash (w/o) feedback 55.28 57 64.03
-(w) feedback 59.37 63 68.98
 Table 2: Performance comparison of VisPath with and without visual feedback.
-The MatPlotBench scores (Plot Score and Executable Rate) and the average score from the Qwen-Agent Code Interpreter benchmark are shown for two LLMs.
-Instead of refining each candidate visualization with feedback from vision-language feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections.
+Results are reported on MatPlotBench (Plot Score, Executable Rate) and the average score on the Qwen-Agent Code Interpreter benchmark for two LLMs.
+Model Setting MatPlotBench Qwen-Agent
+Plot Score Exec. Rate (%) Avg.
+GPT-4o mini w/o visual feedback 63.76 56 58.00
+w/ visual feedback 66.12 60 63.96
+Gemini 2.0 Flash w/o visual feedback 55.28 57 64.03
+w/ visual feedback 59.37 63 68.98
+Instead of refining each candidate visualization with VLM-based feedback, this approach aggregates three candidate codes: each derived from a distinct reasoning path, without intermediate corrections.
 This setup reduces computational overhead and execution time while preserving the benefits of interpretive diversity.
@@
 While full feedback-driven optimization leads to additional performance improvements, this result highlights that the primary strength of VisPath lies in its capacity to explore and leverage diverse reasoning trajectories.
-The framework remains effective and adaptable, even with minimal refinements, further validating importance of its core design.
+The framework remains effective and adaptable, even with minimal refinements, further validating the importance of its core design.
+Table 3: Ablation study isolating the impact of rendered visual feedback. Comparison between structured plot-based feedback and binary execution-only feedback in VisPath.
+LLM Feedback Type Plot Score Exec. Rate (%)
+GPT-4o mini Binary-only feedback 64.82 58
+Visual feedback 66.12 60
+Gemini 2.0 Flash Binary-only feedback 57.68 59
+Visual feedback 59.37 63
 4.3.3 Distinct Contribution of Visual Feedback
@@
 On Gemini 2.0 Flash, Plot Score rises from 57.68 to 59.37 (+1.69), and Executable Rate from 59% to 63% (+4 points).
-Despite the numerical gains are modest, the results demonstrate the unique value of structured visual evaluation.
+Despite the numerical gains being modest, the results demonstrate the unique value of structured visual evaluation.
 Visual feedback enables more refined and user-aligned outputs by capturing subtle rendering issues that may not affect executability, demonstrating its importance in the final synthesis stage.
-LLM Variant Plot Score Executable Rate (%)
-GPT-4o mini (w) feedback 66.12 60
-(w) binary feedback 64.82 58
-Gemini 2.0 Flash (w) feedback 59.37 63
-(w) binary feedback 57.68 59
-Table 3: Ablation results isolating the impact of rendered visual feedback.
-Comparing the full VisPath model using structured plot-based feedback with a binary-only feedback variant.
-Visual feedback provides consistent gains in both Plot Score and Executable Rate across LLMs.
+Figure 3: (a) Scatter plot generation with explicit spatial and annotation constraints.
+User query: “Create a scatter plot of two distinct sets of random data, each containing 150 points.
+The first set (Group X) should be centered around (-2, -2) and visualized in blue, and the second set (Group Y) should be centered around (2, 2) and visualized in orange.
+Label each group at their respective centers with a round white box (…) ”
+(b) Large-scale time-series visualization task.
+User query: “Visualize a large number of time series in three different ways.” This task evaluates each method’s ability to interpret an underspecified request, select appropriate plotting strategies, and compose multiple complementary visualizations for dense temporal data.
+GT denotes the ground-truth visualization.
 5 Discussion
+5.1 Effectiveness of Multi-Path Reasoning and Feedback-Driven Refinement
 Our proposed VisPath framework substantially advances visualization code generation by addressing the core weaknesses of existing methods: limited interpretive flexibility and insufficient refinement.
@@
 Experimental results confirm its superiority: VisPath outperforms all baselines on both MatPlotBench and the Qwen-Agent Code Interpreter benchmark, with up to 9.14% improvement in Plot Score and 10% in Executable Rate.
+Also, Fig.
+3 compares VisPath with the baselines.
+VisPath matches the ground-truth structure more closely when GT is available, and still produces meaningful multi-view visualizations in Fig.
+3(b) where no GT exists.
+The results further show that VisPath correctly follows layout constraints and handles vague time-series prompts, such as centering data groups and creating multiple views for large datasets.
+This suggests that the framework aligns more consistently with user intent in scenarios that remain challenging for other methods.
 Ablation studies further validate VisPath’s design.
@@
 Third, using structured plot-based feedback, rather than binary execution signals, significantly improves output alignment with user intent, confirming the value of our feedback-driven optimization module.
-The figures provided in Appendix clearly illustrate VisPath’s robustness.
-[^4] The detailed cases are provided in Appendix B.
-In Case 1, it correctly centers conditional means using adjusted strip widths, unlike other methods that misalign pointplots or ignore key parameters.
-In Case 2, VisPath generates a proper polar bar chart with radial coordinates and legible labels, while others either fail to render or produce standard rectangular plots, with text often overlapping or unreadable.
-Case 3 highlights VisPath’s ability to handle compositional visualization requests: it correctly interprets the instruction to "visualize in 3 different ways" by generating a multi-subplot layout that includes line, scatter, and bar plots.
-Other approaches, by contrast, either collapsed all time series into a single chaotic plot or failed to differentiate the visualization modalities at all.
-These examples demonstrate how VisPath’s Multi-Path Reasoning and visual feedback lead to more precise and semantically aligned visualizations than current alternatives.
+5.2 Cost & Latency Trade-offs
+To assess the latency impact of Multi-Path Reasoning and structured visual feedback, we compared the iteration counts of VisPath against the MatPlotAgent baseline on MatPlotBench and the Qwen benchmark.
+Iterations were measured across four categories: Query Expansion, Code Generation, Visual Feedback, and Editor.
+While MatPlotAgent uses dynamic iterations based on feedback, VisPath employs a fixed structure of $k=3$ reasoning paths for each query expansion.
+Table 4: Iteration counts across components for MatPlotAgent and VisPath. Benchmarks differ in size: MatPlotBench has 100 rows, while the Qwen benchmark has 80 rows.
+MatPlotBench (100 rows) Qwen (80 rows) MatPlotAgent VisPath MatPlotAgent VisPath Query Expansion 100 100 80 80 Code Generation 312 300 240 240 Visual Feedback 49 300 52 240 Editor 285 100 226 80 Total Iterations 746 800 598 640 Avg Iterations / Row 7.46 8.00 7.475 8.00
+As shown in Table 4, MatPlotAgent averaged 7.46 iterations on MatPlotBench, while VisPath required 8.
+Importantly, VisPath incurs higher iteration counts for Visual Feedback but substantially fewer for the Editor stage (285 vs.
+100 on MatPlotBench; 226 vs.
+80 on the Qwen benchmark).
+Similarly, on the Qwen benchmark, MatPlotAgent averaged 7.475 iterations, with 8 for VisPath.
+This corresponds to a marginal increase of only 0.525–0.54 iterations, despite VisPath delivering substantial improvements in execution success and visual quality.
+The ablation study (2) indicate that accuracy gains do not scale proportionally with $k$, suggesting that efficacy of VisPath is driven by its feedback-oriented design rather than brute-force iteration.
+By redistributing computational effort from editor retries to structured visual feedback, the system achieves meaningful performance without significant overhead.
+Since model calls directly impact latency and cost, $k=3$ provides the effective trade-off between robustness and efficiency for deployment beyond controlled benchmarks.
 6 Conclusion
@@
 Unlike prior methods, our approach seamlessly combines Multi-Path Reasoning with feedback-driven optimization.
-Accurately capturing diverse user intents and refining generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark.
+By accurately capturing diverse user intents and iteratively refining the generated code, VisPath achieves notable improvements in both execution success and visual quality on challenging benchmarks such as MatPlotBench and the Qwen-Agent Code Interpreter Benchmark.
 By prioritizing adaptability, VisPath is uniquely positioned to handle ambiguous user queries through a combination of diverse reasoning paths and visual feedback integration.
 Future work could explore VisPath’s adaptability in more dynamic, real-world scenarios, further broadening its scope and practical utility in complex data analysis contexts.
-7 Limitations
-Despite its effectiveness, the current framework relies on a limited feedback mechanism focused on query-code and query-plot alignment.
-While informative, these signals may overlook finer-grained elements essential to interpretability.
+7 Limitation
+Despite its effectiveness, the current framework focuses on a feedback mechanism that assess query-code and query-plot alignment, which may overlook fine-grained elements essential to interpretability.
 Thus, future work could improve feedback depth by assessing individual plot components, such as readability and visual coherence, enabling more precise and refined visualization code generation.
@@
 Future work could explore ways to selectively identify the most promising reasoning paths early in the process, reducing redundant computation while preserving the benefits of diverse interpretation.
-Appendix A Appendix A. Prompts Used
-Prompt for Multi-Path Reasoning [System Prompt] According to the user query, expand and solidify the query into detailed instruction on how to write python code to fulfill the user query’s requirements.
-Import the appropriate libraries.
-Pinpoint the correct library functions to call and set each parameter in every function call accordingly.
-[User Prompt] Think step by step.
-Generate three distinct extended queries based on the given query.
-Ensure that you first analyze the given data description and create queries that align with the data.
-If no data description is provided, follow the initial query as is.
-You must follow the Python list format for the output.
-Do not modify the detailed instructions from the original user query.
-initial query: oriquery Data description: datadescription Output format: [extendedpath1, extendedpath2, extendedpath3]
-Prompt for Code Generation [System Prompt] You are an expert in data visualization code generation.
-Think step by step and write the generated code in the format “‘python…“‘, where "…" represents the generated code.
-The code must end with ‘plt.show()‘.
-[User Prompt] Think step by step.
-Based on the user’s query and the provided data description, generate Python code using ‘matplotlib.pyplot‘ and ‘seaborn‘ to create the requested plot.
-Ensure that the code is formatted as “‘…“‘, where "…" represents the generated code.
-User query: {query} Data description: {datadescription}
-Prompt for Visual Feedback [System Prompt] Given a code, a user query, and an image of the current plot, please determine whether the plot accurately follows the user query.
-Provide detailed instructions on how to enhance the plot using Python code.
-[User Prompt] Carefully analyze the provided Python code, the user query, and the plot image (if available) to assess whether the generated plot meets the user query requirements.
-If the plot image is missing, check the error message that occurred in the code.
-Compare the plot with the user query, highlight discrepancies, and provide clear, actionable instructions to modify the code.
-Additionally, suggest improvements for better visualization, focusing on clarity, readability, and alignment with the user’s objectives.
-Code: {code} User query: {oriquery}
-Prompt for Synthesis [System Prompt] You are an expert on data visualization code judgement and aggregation.
-[User Prompt] Think step by step.
-Given the provided user query, data description, multiple data visualization codes generated for the query, and feedback for each code’s generated image.
-Your task is to: 1.
-Carefully review the user query and the data description.
-2.
-Examine each version of the data visualization code along with the feedback provided for each version.
-3.
-Synthesize the feedbacks for each code, user query insights, data description to create a final version of the code.
-4.
-Your goal is to produce a final version of code that more effectively fulfills the user query by integrating the best elements from all versions and applying necessary corrections.
-User Query: {oriquery} Data Description: {datadescription} Code for aggregation with corresponding feedback: {codeforaggregation}
-Prompt for Evaluation: MatplotBench You are an excellent judge at evaluating visualization plots between a model generated plot and the ground truth.
-You will be giving scores on how well it matches the ground truth plot.
-The generated plot will be given to you as the first figure.
-If the first figure is blank, that means the code failed to generate a figure.
-Another plot will be given to you as the second figure, which is the desired outcome of the user query, meaning it is the ground truth for you to reference.
-Please compare the two figures head to head and rate them.
-Suppose the second figure has a score of 100, rate the first figure on a scale from 0 to 100.
-Scoring should be carried out in the following aspect: 1.
-Plot correctness: Compare closely between the generated plot and the ground truth, the more resemblance the generated plot has compared to the ground truth, the higher the score.
-The score should be proportionate to the resemblance between the two plots.
-In some rare occurrence, see if the data points are generated randomly according to the query, if so, the generated plot may not perfectly match the ground truth, but it is correct nonetheless.
-Only rate the first figure, the second figure is only for reference.
-If the first figure is blank, that means the code failed to generate a figure.
-Give a score of 0 on the Plot correctness.
-After scoring from the above aspect, please give a final score.
-The final score is preceded by the [FINAL SCORE] token.
-For example [FINAL SCORE]: 40.
-Prompt for Evaluation: Qwen-Agent Code Interpreter benchmark Please judge whether the image is consistent with the [Question] below, if it is consistent then reply "right", if not then reply "wrong".
-Question: {query}
-Appendix B Appendix B. Case Study
-Figure 4: Overview of the end-to-end process of visualizing conditional means alignment.
-Figure 5: Overview of the end-to-end process of generating a polar bar chart.
-Figure 6: Overview of the end-to-end process of visualizing time series data in three different ways.
+Finally, our evaluation primarily focuses on quantitative metrics such as Plot Score and Executable Rate.
+While these metrics offer objective insights, they may not fully capture user-perceived quality, usability, or interpretability of the generated visualizations.
+Conducting user studies or expert reviews could provide complementary qualitative evidence and further validate the practical utility of VisPath.